### v1

# Interpretable Concept Bottlenecks to Align Reinforcement Learning Agents

Abstract
Reward sparsity, difficult credit assignment, and misalignment are only a few of the many issues that make it difficult, if not impossible, for deep reinforcement learning (RL) agents to learn optimal policies. Unfortunately, the black-box nature of deep networks impedes the inclusion of domain experts who could interpret the model and correct wrong behavior. To this end, we introduce Successive Concept Bottlenecks Agents (SCoBots), which make the whole decision pipeline transparent via the integration of consecutive concept bottleneck layers. SCoBots make use of not only relevant object properties but also of relational concepts. Our experimental results provide strong evidence that SCoBots allow domain experts to efficiently understand and regularize their behavior, resulting in potentially better human-aligned RL. In this way, SCoBots enabled us to identify a misalignment problem in the most simple and iconic video game, Pong, and resolve it.
Explaible AI (XAI), Reinforcement Learning, Concept Bottlenecks
1 Introduction
Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffering from a variety of issues that hinder them from learning optimal, generalisable action policies. Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023). A more recent, yet pressing issue is the goal misalignment problem (di Langosco et al., 2022). It occurs when an agent optimizes a different side-goal that is aligned with the target goal at training time, but not at test time (Koch et al., 2021; di Langosco et al., 2022). We discover that such misalignment can occur in the simplest and most iconic ALE (Mnih et al., 2013) game, Pong, in which an agent’s target goal is to play a ball game against an enemy. However, due to the nature of the automated enemy an agent can learn to focus on the position of the enemy paddle for anticipating where to position its own paddle for catching the ball, rather than the position of the ball itself. This can lead to failure of the agent when the enemy is occluded (cf. Fig. 1 (top)). This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020)) in which a model learns good decision rules for the provided training set that, however, fail to transfer to novel data sets. Importantly, if left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings.
Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision. Top: Deep RL agents trained on Pong produce high playing scores ( Refer to caption ) while importance map explanations suggest sensible underlying reasons of taking an action. However, when the enemy is hidden, the deep RL agent fails to even catch the ball ( Refer to caption ) without clear reasons. Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection , e.g., at a relational concept, but also action level. Moreover, they allow for users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts. In this way SCoBots can mitigate RL specific caveats like goal misalignment.
A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI) (e.g. (Schramowski et al., 2020; Lapuschkin et al., 2019)) where the goal of XAI methods is to identify the reasons behind a model’s decisions (Guidotti et al., 2019; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023). Many methods of XAI however suffer from issues such as faithfulness (the explanation does not faithfully present a model’s underlying decision process)(Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020) and low-level semantics (explanations are made at the level of the input data e.g. pixels and not the relevant concepts that are extracted from these pixels)(Stammer et al., 2021; Kambhampati et al., 2021; Teso et al., 2023). A recent branch of research therefore focuses on developing models that provide inherent, concept-based explanations. Prominent examples of this idea are concept bottlenecks models (CBMs) which, among other things, have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020). More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
Despite the success and advantageous properties of CBMs up to now, no research has been performed on their integration into RL, particularly for tackling RL-specific issues discussed above. A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape etc… However, objects in RL environments such as ALE games interact with another (Kaiser et al., 2019). Therefore, in order to adequately solve such tasks additional relational concepts are necessary that, however, build upon these single object concepts.
For this purpose, in this work we introduce Successive Concept Bottleneck Agents (SCoBots). SCoBots (cf. Fig. 1 (bottom)) hereby integrate multiple, successive concept bottlenecks into their decision processes where each bottleneck layer consists of concepts representations that integrate the concept representations of the previous bottleneck layer(s). Importantly, given a set of predefined concept functions that automatically set objects and concepts of initial bottleneck layers in relation to another the final bottleneck layer consists of valuable relational concept representations. SCOBots thus represent inherently explainable RL agents which in comparison to deep RL agents allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes, from single object concepts, to relational concepts to the actual action decision.
In our evaluations on benchmark Atari games we provide experimental evidence not just of the performance abilities of SCoBots that is on par with deep RL agents. More importantly we showcase their ability to provide valuable explanations and their ability to mitigate a variety of RL specific issues, from reward sparsity to misalignment problems. Importantly, in the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents even on the arguably simplest of all Atari games, Pong. Utilizing the advantages of SCoBots this behaviour is, however, easily removed. Ultimately our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
In summary our specific contributions are:
We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their predictions.
We show that SCoBots allow to follow their internal decision processes.
We show that the inherent inspectable nature of SCoBots can be utilized for identifying previously unknown misalignment problems of RL agents.
We show that they allow human interactions that lead to mitigating different RL specific learning issues.
We proceed as follows. We first formally introduce SCoBots and discuss their specific properties. We continue with extensive experimental evaluations on Atari game environments via several analysis and metrics. Before concluding, we touch upon related work.
Figure 2: An overview of Successive Concept Bottlenecks Agents (SCoBots). SCoBots decompose the policy into consecutive interpretable concept bottlenecks (ICB). Objects and their properties are first extracted from the raw input, human-understandable functions are then employed to derive relational concepts and finally used to select an action. The understandable concepts enable interactivity. Each bottleneck allows expert users to, e.g., prune or utilize concepts to define additional reward signals.
2 Successive Concept Bottleneck Agents
Let us start with providing the background notations on RL via markov decision processes (MDPs) as well as concept bottleneck models (CBMs). In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment. Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ with parameters $\theta$, are usually black-box models that process raw input states, e.g. a set of frames, and provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
Concept bottleneck models (CBMs) initiate a learning process by extracting relevant concepts from raw input e.g. image data. These concepts can e.g. represent the color or position of a an Atari character. Subsequently, these extracted concepts are used for downstream tasks such as classification tasks. Formally, a bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into a concept representation c. Next, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$.
The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects. In contrast, RL agents must learn not just from static, labeled data but also through interaction with dynamic environments. RL tasks thus often necessitate relational reasoning, as they involve understanding the relationships between depicted instances that evolve through time and interact with one another. This is crucial for learning effective policies in complex, dynamic environments. In this manuscript, we use specific fonts to distinguish object properties and relations.
2.1 Building inspectable ScoBots
An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user. This stands in contrast e.g. to unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020) where there is no guarantee for learning human-relatable concepts. SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations. These steps are described herafter and depicted in Fig. 2.
Similar to other RL agents, SCoBots process the last $n$ frames $\{x_{t}\}_{t=T-n}^{T}$, observed from the environment. For each frame, SCoBots employ an initial concept extraction method as is standard in other works on CBMs (Stammer et al., 2021; Koh et al., 2020) to extract objects and their properties. Specifically, we start from an initial bottleneck model, $\omega_{\theta_{1}}(\cdot)$, that is parameterized by parameter set, $\theta_{1}$. Given an image, $x_{t}$, this model provides a set of $c_{t}$ object representations, $\omega_{\theta_{1}}(x_{t})=\Omega_{t}=\{o_{t}^{j}\}_{j=1}^{c_{t}}$, where each object representation corresponds to a tensor of different extracted properties of that object, (e.g. the object category, position (i.e. its $x,y$ coordinates), size, … etc). As in previous works on CBMs this bottleneck model, $\omega_{\theta_{1}}(\cdot)$, can correspond to a model that was supervisedly pretrained for predicting these object properties from an image.
One of the major differences to previous work on CBMs is that SCoBots now extract and utilize relational concepts that are based on the previously extracted objects and their properties. Formally, SCoBots utilize a successive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor. This model, $\mu$, is parameterized by a predetermined set of relational functions, $\mathcal{F}$, that are used to extract a set of $d_{t}$ relational concepts. These relations are based on each individual object or property for unary relations or combinations of these for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$. Without strong prior knowledge $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed. However, this set can easily be updated on the fly by a human user e.g. with additional, novel relation functions.
Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (e.g. image $x_{t}$) and that is parameterized by $\theta_{2}$}. Up to now all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations. To guarantee understandability also within the action selection step of SCoBots we require to embed a specifically interpretable action selector. While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods they lack this required interpretability feature. However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models. Thus, to trade-off the issues of flexibility and performance vs interpretability SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization. After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$. Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
Overall, our approach preserves the MDP formulation used by classic deep approaches, but decomposes the classic deep policy (1) into a successive concept bottleneck one (2):
$s_{t}$ $\xrightarrow{\pi_{\theta}}a_{t}$ (1)
$s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}$ $\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$ (2)
where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) are the policy parameters. For simplicity, we will discard the parameter notations in the rest of the manuscript. Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in App.A.6 and App. A.7.
Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently. This separation of training for different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem. In this way, the object-concept extractor, $\omega$, can be optimized by supervised (Redmon et al., 2016) or unsupervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods. On the other hand, the set of suitable relations, $\mathcal{F}$, as well which objects to apply them on, may be adapted via evolutionary algorithms (Peng et al., 2018; Yang et al., 2022), supervision from large language models (Wu et al., 2023), via prior task-knowledge (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b). Conclusively, the action selector, $\rho$ can be optimized based on the direct RL signal.
2.2 Guiding SCoBots
The inspectable nature of SCoBots not only brings the benefit of improved human-understandability, but importantly allows for targeted human-machine interactions. In the following we describe two guidance approaches for interacting with the latent representations of SCoBots: concept pruning and object-centric rewarding.
Concept pruning. The type and amount of concepts that are required for a successful policy may vary greatly across tasks. For instance, object colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman. At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g. the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$). Moreover, for a single RL task, the need for specific concepts might even change during training, e.g. in progressive environments where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2021).
Specifically, SCoBot’s human-comprehensible concepts allow domain experts to prune unnecessary concepts. In this way expert users can guide the learning process towards relevant concepts. Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$. Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, and (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$. iSCoBots thus formally refine Eq. 2, leading to:
[…]
Furthermore, users can (iv) prune out redundant actions resulting in a new action space $\overline{\mathcal{A}}$.
Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections. Human-normalized scores (red line) of deep and object-centric agents trained using PPO on 99 ALE environments. SCoBots obtain similar or better performance than the deep agents, showing object-centric agents can also learn to solve RL tasks while making use of human-understandable concepts (left). Guiding SCoBots allow to correct misalignment in Pong (center) and the ill-defined reward signal of Kangaroo (right).
To give brief examples of these four pruning possibilities we refer to Fig. 2, focusing here on the orange subsections. Particularly, in Pong a user can remove objects (scores) or specific object properties (R,G,B values) which would result in a $\overline{\Omega_{t}}$. Second, the color relation, color($\cdot$), is not relevant for successfully playing Pong and can therefore be removed entirely from $\mathcal{F}$. Third, a user can state that the vertical distance function (dist($\cdot,\cdot$).y) should not be applied to the (player, enemy) input couple. These pruning actions result in the set of relational concepts $\overline{\Gamma_{t}}$. Lastly, the only playable actions in Pong are UP and DOWN. To ease the search, the available FIRE, LEFT and RIGHT from the base Atari action space might be discarded as they are equivalent to NOOP (i.e. no action), resulting in $\overline{\mathcal{A}}$.
Importantly, being able to prune concepts can help mitigate misalignement issues such as those of agents playing Pong (cf. Figure 1), where an agent can base it’s action selection on the enemy’s position. Specifically, by pruning the enemy position from the distance relation concept a user can guide the SCoBot to rely on relevant features for their decisions such as the ball’s position, rather than the spurious correlation with the enemy’s paddle.
Object-centric feedback reward. Reward shaping (Touzet, 1997; Ng et al., 1999) is a standard RL technique that is used to facilitate the agent’s learning process by introducing intermediate reward signals, aggregated to the original reward signal. The object-centric and importantly concept-based approach of SCoBots allows to easily craft additional reward signals. Formally, one can use the extracted relational concepts to express new expert reward signal:
$R^{\text{exp}}(\Gamma_{t})$ $:=\sum_{g_{t}\in\Gamma_{t}}\alpha_{g_{t}}\cdot g_{t},$ (4)
where $R^{\text{exp}}:\Gamma\longrightarrow\mathbb{R}$. Moreover, where $\Gamma_{t}=\mu(\omega(s_{t}))$ is the relational state, extracted by the relation extractor of a SCoBot. The coefficient $\alpha_{g_{t}}\in\mathbb{R}$ is used to penalize or reward the agent proportionally to relational concepts. Our expert reward only relies on the state, as we make use of the concepts extracted from it. However, incorporating the action into the reward computation is straightforward. In practice, this expert reward signal can lead to guiding the agent towards focusing on relevant concepts, but also help smoothing sparse reward signals, which we will discuss further in our evaluations.
For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf. red components of Fig. 2), with the intention of incentivizing the agent to maintain close proximity with the ball. As we show in our experimental evaluation, we can use this concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf. App. A.9 for details on each problem).
3 Experimental Evalautions
In our evaluations, we investigate several properties and potential benefits of SCoBots. We specifically aim to answer the following research questions:
  • Q1. Are SCoBot agents able to learn competitive policies on benchmark RL environments?
  • Q2. Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes, e.g. policy misalignments?
  • Q3. Can concept-based guidance let us mitigate RL caveats of SCoBot agents, e.g. policy misalignments?
Figure 4: SCoBots allow to follow their decision process. Their object centricity and interpretable concepts allow us to follow their reasoning. The decision processes of SCoBots (extracted from the decision trees) are provided with their associated states for an guided SCoBot on Skiing (left) and a SCoBot on Pong (right). For examble, in this Skiing state, our iSCoBot selects RIGHT, as the distance between Player and the (left) Flag1 (on the $x$ axis) is high (bigger than $15$). This agent selects the correct action for the right reasons.
Experimental setup: We evaluate SCoBots on $9$ Atari games (v5 version) from the ALE benchmark (Bellemare et al., 2012) (by far the most used RL framework (cf. App. A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game. For results based on achieved game scores we provide human normalized scores (following eq. 5) that are averaged over $3$ seeds for each agent configuration. We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and specify these as neural agents in the following. Note that we evaluate these agents on the latest v5 version of the environments, leading to slightly worst results than (Schulman et al., 2017). We obtain similar results when evaluating on the old versions (cf. Tab. 1). For the SCoBots we assume access to a pretrained object extractor and thus provide our agents object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a). This allows us to focus our evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g. (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)). For all agents, we restrict training to $20\text{M}$ frames and use the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 (Raffin et al., 2021) implementation and its default hyperparameters for both the deep neural and SCoBot agents (cf. Tab. 2 in the Appendix). SCoBots’ inputs consist of vectors containing the values of concatenated concepts (i.e. properties and relations). Specifically, in our evaluations, the set of object properties consists of object class (e.g. enemy paddle, ball etc.), $(x^{t},y^{t})$ coordinates, coordinates at the previous position $(x^{t-1},y^{t-1})$, height and width, the most common RGB values (i.e. the most representative color of an object), and object orientation. The specific set of functions, $\mathcal{F}$, that are used to extract object relations is composed of: euclidean distance, directed distances (on $x$ and $y$ axes), speed, velocity, plan intersection, the center (of the shortest line between two objects), and color name (cf. App. A.6.1 for implementation details). To distill SCoBots learned policy of SCoBots, we use the rule-based model extraction method from (Zarlenga et al., 2021). For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf. Tab. 4). Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below. More details about the experimental setup and the hyperparameters are provided in App. A.6.
SCoBots learn competitive policies (Q1). We present human-normalized (HN) scores of both SCoBot and neural agents trained on each investigated game individually in Fig. 3 (left). Exact numerical results are provided in Table 1 (cf. App. A.3). Overall, our results suggest that SCoBots provide competitive performances on every tested game despite the constraints of multiple bottleneck within their architectures. This shows algorithms where the object extraction is decoupled from the reasoning process that leads to the action selection can be developed. Other environments might require additional information extraction processes (e.g. MsPacman, in which the agent must navigate a maze), and we haven’t evaluated on such environments here. This constitute a limit to our SCoBots.
Inspectable SCoBots’ to detect misalignments (Q2). A main target of developing ScoBots is to obtain competitive, yet transparent RL agents that allow for human users to identify an agent’s underlying reasoning process. Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allow to pinpoint not just the object properties, but importantly also the relational concepts and properties that are being used for the final action decision. This is exemplified in Fig. 4 for the games of Skiing and Pong (cf. App. A.5 for explanations of SCoBots on the remaining games). Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of the two games. For example, for the specific game state in Skiing, the SCoBot agent selects the action RIGHT as the best option, because the distance from its character to the left flag is larger than a specific value ($15$ pixels). Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
A more striking display of the benefits the inherent explanations of SCoBot agents can be found in the Pong-based agent in Fig 4. Specifically, the provided explanation suggests that the agent is basing its decision on the vertical position of the enemy’s and of its own paddle (distance between the two paddles y direction). This, in turn, suggests that in fact the agent is ignoring the ball’s position, at least to a certain degree. Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay). Thus, its vertical position is highly correlated with the enemy’s one (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf. App. A.8). Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.
To validate these findings further we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (we refer to this environment as NoEnemy cf. Fig. 1). We evaluate both the neural and SCoBot agents that were trained on the original Pong environment and observe catastrophic performance drops in HN scores (cf. Fig. 3 both of SCoBot and both neural PPO) at a level of random policies, when evaluated on this environment (NoEnemy). We have also run an DQN agent on this environments and observed the same results. The drop in neural agent performance is particularly striking as initial importance maps produced by these neural agents had highlighted all $3$ moving objects as relevant for their decisions (i.e. the player, ball, and enemy) (cf. Fig.1 top left and App. A.8, aligned with findings of (Weitkamp et al., 2018)). These importance maps suggest the agents had based their decision on right reasons. Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and but more importantly the granularity of such post-hoc explainability methods in the context of RL. It highlights the importance of inherently explainable RL models on the level of concepts as provided via SCoBots to identify potential issues such as goal misalignment. In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
Guiding SCoBots to mitigate RL-specific caveats (Q3).
We here make use of SCoBots interaction capabilities to address $4$ famous RL specific problems: goal misalignment which we had previously identified via SCoBot’s inherent, concept-based explanations and also sparse rewards, ill-defined rewards and difficult credit assignment.
Realigning SCoBots: For mitigating identified SCoBots goal misalignment issues, we can correct the previously introduced misalignment, by simply removing (pruning) the enemy in the set of considered objects ($\overline{\Omega}$). Doing so, the enemy cannot be used for the action selection process. This leads to an agent that is able to play Pong and its NoEnemy version (guided SCoBot in Fig 3 (center)). This shows again that playing Pong without observing the enemy is achievable, and thus should also be the case for deep agents.
Sparse Reward RL: Let us move on to the issue of sparse reward in the context of RL. In Pong, to score a point, the player needs to return the ball and have it go past the enemy. To do so, the player has to perform a spiky shot, obtained by touching the ball on one of its paddle’s sides (otherwise, the shot is flat and the enemy is easily catching it). The probability of such a shot is approximately $3.8\%$ (as shown in details in Appendix A.2). Thus, a randomly playing (or untrained) agent is only rewarded on average every $2200$ steps. Providing a SCoBot agent with an additional reward that is inversely proportional to the distance between its paddle and the ball incentivizes the agent to keep a vertical position close to the ball’s one, tripling the probability of a winning shot in our experiment (cf. App. A.2). Thus, the reward signal in Pong is low, but our example illustrates how the interpretable concepts allow to easily guide SCoBots, making up for the reward sparsity.
Ill-defined reward: As previously explained, defining a reward upfront that incentivizes agents to learn the expected behavior can be difficult. An example of ill-defined reward (i.e. a reward that will lead to an unaligned behavior if the agent maximizes the return) is present in Kangaroo. According to the documentation of the game111www.retrogames.cz/play_195-Atari2600.php, “the mother kangaroo on the bottom floor tries to reach the top floor where her joey is being held captive by some monkeys”. However, punching the monkey enemies gives a higher cumulative reward than climbing up to save the baby. RL agents thus tend to learn to kick monkeys on the bottom floor and never go to the joey (cf. Fig. 4). For revising such agents (i.e. incentivize them to reach the games’ original intended goal), we provide the SCoBot agents with an additional reward signal, based on the agent’s progression to the joey, (cf. App. A.9). As can be seen in Fig. 3 this reward allows us to obtain SCoBot agents that achieve high HN game scores Fig. 3 (left) that indeed complete the level Fig. 3 (right). In contrast the neural agents achieve high HN scores, but do not complete the levels, whereas unguided SCoBot agents achieve lower HN scores, but higher level completion ratios than the neural agents, but greatly lower ones than the guided SCoBot agents.
Difficult Credit Assignment Problem: Lastly, the difficult credit assignment problem in RL is the challenge of correctly attributing credit or blame to past actions when the agent’s actions can have delayed and non-obvious effects on its overall performance. We illustrate this in the context of Skiing, where in standard configurations agents receive at each step a negative reward that corresponds to the number of seconds elapsed since the last steps (i.e. varying between 3-3 and 7-7). This reward aims at punishing agents for going down the slope slowly. Additionally, the game keeps track of the number of flags that the agent has correctly traversed (displayed at the top of the screen, cf. Fig. 4). A large positive reward, proportional to it, is attributed to the agent when it reaches the final state. Associating this large reward signal with the passed flags is extremely difficult, without prior knowledge on human skiing competitions. At each timestep, in the next evaluations we thus provide SCoBot agents with a reward that is proportional to the agent’s progression to the flags’ position to incentivize an agent to pass in between them as:
$R^{\text{exp}}_{1}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}-D(Player,o)^{t-1}.$
This resulted in an agent that learned to stop itself in between the flags (depicted in Fig. 12), showing once again, how difficult it is to create a reward that would favor a specific behavior. To correct it, we therefore adjoined another signal to reward our SCoBot agents proportionally to the speed of the agent:
$R^{\text{exp}}_{2}=V(Player).y.$
These rewards are then summed to the environmental one. With this we obtain guided SCoBot agents that reach human scores, whereas the neural and non-interactive SCoBot agents perform worse than randomly (cf. Fig. 3 (left)).
Overall, our experimental evaluations show that the not only are SCoBots on par with neural RL agents, but that importantly the concept-based and inspectable nature of SCoBots allows identify and revise several important RL specific caveats. In the course of our evaluations we could identify several previously unknown issues of RL agents on benchmark environments even as simple as Pong.
4 Related Work
The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as (Lampert et al., 2009) and (Kumar et al., 2009). A first systematic study of CBMs and their variations was delivered by (Koh et al., 2020) and (Stammer et al., 2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output. Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$. Since the initial works on CBMs they have found utilization in several applications. (Antognini & Faltings, 2021) e.g. apply CBMs to text sentiment classification. (Bahadori & Heckerman, 2020) propose a bias-removing method to tackle the problem of confounding concept correlations. (Chauhan et al., 2022) on the other hand query a domain expert for concept adaptation at prediction time. However, all of these works consider only single object concept representations and mainly focus on supervised learning settings.
In fact, CBMs have found their way into RL only to a limited extent. (Zabounidis et al., 2023) and (Grupen et al., 2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance. (Zabounidis et al., 2023) further report better training stability and reduced sample complexity. Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part. SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations. Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, (Guan et al., 2023) allow the user to design a reward function based on “behavioral attributes”, i.e. higher-level properties that occur not in a single state but over a period of time. Lastly, SCoBots not only separate state representation learning from policy search, as done by (Cuccu et al., 2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022). For example, the true qualities of a good explanation are unclear and there is a lack of comparative explainability evaluation benchmarks. (Milani et al., 2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior. Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods. According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining). Moreover, (Dazeley et al., 2021) suggest different “levels of explanation” based on the agent’s intentions. SCoBots especially facilitate the generation of zero-order explanations, i.e., reactive explanations that explain the agent’s reaction of an immediately perceived input. Overall, SCoBots transform the raw state $s$ into a symbolic state $s^{\prime}=\mu(\omega(s))$ in line with the proposal of Kambhampati et al.. d’Avila Garcez et al. do an analogous symbolic extraction but resulting in a set of symbolic “sub-states” $s_{k}^{\prime}$ instead of $s^{\prime}$: Each $s_{k}^{\prime}$ represents the relative position of the agent to one interactable object. This way, however, solving one MDP results in solving a set of partially observable MDPs which is avoided in the case of SCoBots.
The misalignment problem is related to shortcut learning, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023) to prototype-based models (Bontempelli et al., 2023) to reinforcement learning (RL) approaches (di Langosco et al., 2022). A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022). The authors (Nahian et al., 2021) align the agent to ethical values by introducing a second reward signal. In comparison, SCoBots aid to resolve RL specific learning issues such as the misalignment problem through inherent interpretability.
5 Conclusion
In this work, we have shown the benefits of concept-based models in RL tasks, specifically for identifying issues such as goal misalignment. Among other things our proposed Successive Concept Bottleneck agents specifically integrate automatically generated relational concept representations into their decision processes. With this we could identify previously unknown misalignment problems of deep RL agents in games as simple as Pong. Due to their inspectable and human-understandable nature of our proposed SCoBot agents these issues could not only be identified, but easily revised with minimal additional feedback. Our work thus represents an important step in developing aligned RL agents, i.e. agents that are not just aligned with the underlying task goals, but also with human user’s understanding and knowledge. Achieving this is particularly valuable for applying RL agents in real-world settings where ethical and safety considerations are paramount.
Avenues for future research are incorporating options, which represent higher-level actions as described in (Bacon et al., 2017), in order to make our (i)SCoBots more closely resemble human-like decision-making processes. One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023). Additionally, we are considering the use of shallower decision trees, as proposed by (Broelemann & Kasneci, 2019), to enhance interpretability further in our RL agents. In our experiments we assume the initial object concept representations have been extracted. An interesting research question is further in how far the task reward signal can aid in learning such object-centric representations in the first place e.g. based on the MOC framework of Delfosse et al..
Appendix A Appendix
As mentioned in the main body, the appendix contains additional materials and supporting information for the following aspects: further information on the fact that Atari is the most common set of games (A.1), details on the reward sparsity in Pong (A.2), detailed numerical results (A.3), learning curves of our RL agents (A.4), the hyperparmeters used in this work (A.6), formal definitions on the SCoBots policies (A.7) and further details on the misalignment problem in Pong.
A.1 Atari games are most common set of environments
We here show that the Atari games from the Arcade Learning Environment (Bellemare et al., 2012) is the most used benchmark to test reinforcement learning agents.
Figure 5: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from (Delfosse et al., 2023a).
A.2 The reward sparsity of Pong.
In the game Pong, the agent has a low chance of discovering positive reward. Its vertical position may vary between $34$ and $190$. Its paddle height is $16$, but it needs to shoot from the side of the paddle (the $3$ pixel of each border) to get a spiky shot that is likely to go beyond the opponent. If the balls arrives not close to the top or bottom borders, the enemy will still catch it, which has the probability of $\sim 22\%$ in our experiments. The probability of getting a successful shot is thus of $\sim 6/156*0.78=3\%$. The average number of corresponding tryouts for the agent to get rewarded is $\sum_{n=1}^{\infty}\left(n\times 0.03\times 0.97^{n-1}\right)=33.3$. The agent usually need $60$ steps (in average) from the initialization of the point to the reward attribution, hence, it will need $\sim 2200$ steps to be rewarded. This is consistent with the experiment depicted in Figure 6. In this figure, a random agent (original) needs in average $2230$ steps to observe reward.
Figure 6: Expert user feedback allow for faster discovering of the reward signal in the sparse reward Pong environment.
By providing the agent with a punishment proportional to the vertical distance between its paddle and the ball, we were able to obtain a winning shot every $\sim\!820$, multiplying by $2.7$ their occurrences.
A.3 Detailed numerical results
For completeness, we here provide the numerical results of the performances obtained by each agent type. For fair comparison, we reimplemeneted the deep PPO agents, and used the default hyperparameters for Atari from stable-baselines3 for the deep RL agents. A detailed overview of hyperparameters and their corresponding values can be found in A.6.
All agents are trained on gymnasium’s atari learning environments using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank). Each training seed’s performance is evaluated every 500500k frames on $4$ differently seeded (42+42+training seed) environments for $8$ episodes each. After training, the best performing checkpoint is then ultimately evaluated on $4$ seeded ($123,456,789,1011$) test environments. The final return is determined by averaging the return over $5$ episodes per training-test seed pair and every training seed for the respective environment. We use deterministic actions for both evaluation and testing stages.
Game SCoBots SCoBots (guided) Neural(v5) Original reproduced (v4) Original Random Human
Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433.3\mbox{\scriptsize$\pm 2772.9$}$ $4532$ $210$ $8503$
Bowling $106.6\mbox{\scriptsize$\pm 41.8$}$ $137.4\mbox{\scriptsize$\pm 24.1$}$ $102.2\mbox{\scriptsize$\pm 39.1$}$ $51.4\mbox{\scriptsize$\pm 18.0$}$ $40.1$ $23.1$ $160.7$
Boxing $97.1\mbox{\scriptsize$\pm 2.2$}$ $69.0\mbox{\scriptsize$\pm 10.9$}$ $0.3\mbox{\scriptsize$\pm 3.0$} 99$$9.5\mbox{\scriptsize$\pm 1.4$}$ $94.6$ $0.10$ $4.3$
Freeway $32.8\mbox{\scriptsize$\pm 0.1$}$ $2.9\mbox{\scriptsize$\pm 0.7$} 33$$6\mbox{\scriptsize$\pm 0.2$} 33.$$3.6\mbox{\scriptsize$\pm 0.3$}$ $32.5$ $0.00$ $29.6$
Kangaroo $2776.6\mbox{\scriptsize$\pm 1332.4$}$ $4050.0\mbox{\scriptsize$\pm 217.8$}$ $790.0\mbox{\scriptsize$\pm 280.8$}$ $13296.7\mbox{\scriptsize$\pm 1111.1$}$ $9929$ $52.0$ $3035$
Pong $17.2\mbox{\scriptsize$\pm 1.9$}$ $17.5\mbox{\scriptsize$\pm 1.8$}$ $16.4\mbox{\scriptsize$\pm 1.5$}$ $20.9\mbox{\scriptsize$\pm 0.2$}$ $20.7$ 20.7-20.7 $14.6$
Seaquest $1055.3\mbox{\scriptsize$\pm 272.6$}$ $2411.3\mbox{\scriptsize$\pm 377.0$}$ $837.3\mbox{\scriptsize$\pm 46.7$}$ $1262.0\mbox{\scriptsize$\pm 446.1$}$ $1204.5$ $68.4$ $20182$
Skiing $23004\mbox{\scriptsize$\pm 10333$}$ $6530\mbox{\scriptsize$\pm 3326$}$ $3004\mbox{\scriptsize$\pm 10333$} 22$$2983\mbox{\scriptsize$\pm 10333$}$ 13901-13901 17098-17098 4336-4336
Tennis $2.4\mbox{\scriptsize$\pm 3.6$}$ $0.0\mbox{\scriptsize$\pm 0.0$}$ $0.9\mbox{\scriptsize$\pm 0.1$}$ $1.2\mbox{\scriptsize$\pm 0.3$}$ 14.8-14.8 23.8-23.8 8.3-8.3
Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs. Deep refers to the agent that use a CNN instead of our ICB layers. We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)). Our agents have been trained with only 2020M frames.
Normalisation techniques. To compute human normalised scores, we used the following equation:
[…]
A.4 Learning curves
Figure 7: Training Return over frames ($\times 10^{6}$) seen for   guided SCoBots,   SCoBots and   neural agents.
A.5 Agent reasoning through the decision trees.
In this section, we provide more decision trees from SCoBots on states from Freeway and Bowling.
Figure 8: Decision trees from Freeway and Bowling.
A.6 Hyperparameters and experimental details
Actors NN 8
Minibatch size 32832*8
Horizon $T$ 2048
Num. epochs $K$ 33
Adam stepsize $2.5*10^{-4}*\alpha$
Discount $\gamma$ $0.99$
GAE parameter $\lambda$ $0.95$
Clipping parameter $\epsilon$ $0.1*\alpha$
VF coefficient $c_{1}$ 11
Entropy coefficient $c_{2}$ $0.01$
Table 2: PPO Hyperparameter Values. $\alpha$ linearly decreases from 11 to 0 over the course of training.
All Experiments were run on a AMD Ryzen 7 processor, 64GB of RAM and one NVIDIA GeForce RTX 2080 Ti GPU. Training a (i)SCoBots on $20\text{M}$ frames and using $8$ actors takes approximately $4$ hours. We use the same PPO hyperparameters as the Schulman et al. (2017) agents that learned to master the games. For the Adam optimizer, SCoBots start with a slightly increased learning rate of $1\times 10^{-3}$ (compared to $2.5\times 10^{-4}$). The PPO implementation used and the respective MLP hyperparameters are based on stable-baselines3 Raffin et al. (2021). SCoBots have the same PPO hyperparameter values as deep agents but use MLPs ($2\times 64$) with ReLU activation functions as policy and value networks. Deep agents use CnnPolicy in stable-baselines3 as their policy value network architecture, which aligns with the initial baseline implementation of PPO. Further, we normalize the advantages in our experiments, since it showed only beneficial effects on learning. This is the default setting in the stable-baseline3 implementation.
The Atari environment version used in gymnasium is NoFrameskip-v4 for agents reproducing the reported PPO results, and v5 for SCoBots and neural. v5 defines a deterministic skipping of 5 frames per action taken and sets the probability to repeat the last action taken to 0.25. This is aligned with recommended best practices by Machado et al. (2018). The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation222github.com/openai/baselines. This includes frame skipping of 4 and reward clipping. Frame stacking is not used. SCoBots are not trained on a clipped reward signal. For comparability, the neural agents we compare SCoBots to, are not as well. A list of all hyperparameter values used is provided in Table 2.
A.6.1 The properties and features used for SCoBots
In this paper, we used different properties and functions (to create features). They are listed hereafter. Further, Table 4 shows a detailed overview of the concepts and actions pruned for every environment evaluated.
Name Definition Description
Properties class NAME object class (e.g. ”Agent”, ”Ball”, ”Ghost”)
position $x,y$ position on the screen
position history $x_{t},y_{t},x_{t-1},y_{t-1}$ position and past position on the screen
orientation $o$ object’s orientation if available
RGB $R,G,B$ RGB values
Relations distance $D_{x}(o_{1},o_{2}),D_{y}(o_{1},o_{2})$ distance the $x$ and $y$ axis
euclidean distance $D_{e}(o_{1},o_{2})$ euclidean distance
trajectory $landing\_point_{x}(o_{1},o_{2})$ orthogonal projection of $o_{1}$ onto $o_{2}$ trajectory
center $center(o_{1},o_{2})$ center of the shortest line between $o_{1}$ and $o_{2}$
speed $s(x_{t},y_{t},x_{t-1},y_{t-1})$ speed of object (pixel per step)
velocity $v(x_{t},y_{t},x_{t-1},y_{t-1})$ velocity of object (vector)
color $color(o_{1})$ the CSS2.1333https://www.w3.org/TR/CSS2/ color category (e.g. Red)
top k $Top_{k}((o_{1},...,o_{n}),k,class)$ only $k$ closest ($D_{e}$ to player) objects visible
Table 3: Descriptions of properties and relations used by SCoBots.
Features Bowling Boxing Pong Freeway Tennis Skiing Kangaroo Asterix Seaquest
Properties class no enemy
position
position history X
orientation X X X X X X X
RGB X X X X X X X X X
Relations trajectory X X X X X X X X
distance X
euclidean distance X X X X X X X X X
center X X X X X X X X
speed X X X X X X X
velocity X X X X X
color X X X X X X X X X
k closest objects 4 X X 4 X 2 2 X X
Actions NOOP
FIRE - - -
UP - -
RIGHT - -
LEFT - -
DOWN - -
UPRIGHT - - - - -
UPLEFT - - - - -
DOWNRIGHT - - - - X -
DOWNLEFT - - - - X -
UPFIRE - - - X - -
RIGHTFIRE - X - - X - -
LEFTFIRE - X - - X - -
DOWNFIRE - - - X - -
UPRIGHTFIRE - - - X - X - -
UPLEFTFIRE - - - X - X - -
DOWNRIGHTFIRE - - - X - X - -
DOWNLEFTFIRE - - - X - X - -
Table 4: Feature selection and pruning for guided SCoBots. denotes the included features, whereas X the features that are pruned out.
A.7 SCoBots policies: formal definitions
$\mathcal{S}$$\mathcal{O}\mathcal{F}\mathcal{R}$$\mathcal{A}$$\omega$$\mu$$\rho$
Figure 9: Functional summary of the SCoBots model architecture: $\omega$ maps states to sets of objects, $\mu$ maps sets of objects to relation vectors by applying relational functions, and $\rho$ maps relation vectors to actions.
The set $\mathcal{S}$ denotes the problem-specific set of raw environment states as defined by the RL problem, e.g., the space of RGB images $0,255]^{512\times 512\times 3}$. Similarly, $\mathcal{A}$ represents the action space, e.g., “move right” or “jump.”
Let $O_{0}$ be the set of all possible objects, identified by their ID and characterized by their properties like position, size, color, etc. Then, $\mathcal{O}:=\{O\in\mathcal{P}(O_{0})\mid\text{id}(o_{1})\neq\text{id}(o_{2})\forall o_{1},o_{2}\in O\}$ is the object detection space, defined as the family of object sets in which each object (identified by its ID) occurs once at most.
$\mathcal{F}$ is the relation function space. It is the family of user-defined relation function sets $F$. Each relation function $f\in F$ is of the form $f:O^{k}\rightarrow\mathds{R}^{n}$ for $k,n\in\mathds{N}$, that is, it maps a fixed number of objects to a real-valued vector. As an instance, the function that maps two objects to the Euclidean distance between each other is a relation function.
The set $\mathcal{R}\subseteq\mathds{R}^{m}$ is the relation (or feature) space. The dimension $m$ is implied by $\mathcal{F}$. More specifically, $m=\sum_{f\in\mathcal{F}}\dim(\text{co}(f))$, i.e., $m$ is the sum of each relation function’s codomain dimension.
Definition A.1.
The overall model policy $\pi:\mathcal{S}\rightarrow\mathcal{A}$ given relation function set $F$ is defined as $\pi:=\rho\circ\mu(\cdot,F)\circ\omega$, where
  1. 1.
    $\omega:\mathcal{S}\rightarrow\mathcal{O}$ is the object detector, defined as the function that maps a raw state $s\in\mathcal{S}$ to a set of detected objects $O\in\mathcal{O}$.
  2. 2.
    $\mu$ is the feature selector
    $\mu:\mathcal{O}\times\mathcal{F}$ $\rightarrow\mathcal{R}$ (6)
    $(O,F)$ $\mapsto\mathbf{r}:=\left(f(o_{1},...,o_{k})\right)_{f\in F},\text{where }o_{1},...,o_{k}\in O.$ (7)
    That is, $\mu$ applies each relation function $f\in\mathcal{F}$ to the respective detected object(s) in $O\in\mathcal{O}$, resulting in a real-valued relation vector.
  3. 3.
    $\rho:\mathcal{R}\rightarrow\mathcal{A}$ is the action selector that assigns actions to relation vectors.
See also Figure 9 for a summary. The policy parameters $\theta:=(\theta_{1},\theta_{2},\theta_{3})$ split up into the parameters of the object detector, the feature selector, and the action selector, accordingly. They are left out for brevity.
The architecture’s bottleneck is induced by the object detector $\omega$ and the feature selector $\mu$. From this function perspective, the bottleneck consists of two stages: an object-centric bottleneck stage at $\mathcal{O}$ and a semantic bottleneck stage at $\mathcal{R}$.
A.8 Pong misalignment problem.
The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019). Here, an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time. When the coin is repositioned at test time di Langosco et al. observed that trained deep RL agents avoid the coin and simply target the end of the level, indicating that the agents have in fact learned to follow a simpler side-goal rather than the underlying strategy.
To find the correlation between the enemy’s and the ball’s y positions, we let an random agent play the game for $100000$ frames and collect the positions of the enemy and ball every $10$ frames. We then compute different correlation coefficients, and obtain 99.6% as Pearson coefficient, 96.4% as Kendall coefficient and 99.5% as Spearman coefficient.
Here are then importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2021).
Figure 10: Importance maps of trained deep DQN agents playing Pong (left) show that all $3$ moving objects are important for the agent’s decision, suggesting that the agent takes aims at sending the ball past the enemy. The same trained agent completely outperform its enemy in the original environment while it cannot return the ball in a modified version of the game where the enemy is hidden.
Figure 11: More importance maps of DQN agents playing Pong, with ReLU (left) and rational activation functions (right).
A.9 RL specific caveats
In this section, we give more details on the different RL specific caveats, as well
Sparse Reward: Some RL environments provide the agent with a rarely occurring reward signal, making it challenging to learn effectively. Such sparse reward environments are particularly present in the robotic field (Andrychowicz et al., 2017).
Ill-defined Objectives: Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018). Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common (Andrychowicz et al., 2017).
To realign our SCoBots agent, we provide them with a reward signal proportional to the progression to the joey:
1player = _get_game_objects_by_category(game_objects, ["Player"])
2
3# Get current platform
4platform = np.ceil((player.xy[1] - player.h - 16) / 48) # 0: topmost, 3: lowest platform
5
6# Encourage moving to the child
7if not episode_starts and not last_crashed:
8 if platform % 2 == 0: # even platform, encourage left movement
9 reward = - player.dx
10 else: # encourage right movement
11 reward = player.dx
12
13 # Encourage upward movement
14 reward -= player.dy / 5
15else:
16 reward = 0
Difficult Credit Assignment: When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021). This problem arises in games like chess, where reward is passed to the agent only when the game is over. The problem can be addressed by more instant informative feedback, given directly after successful actions or after a positive state is reached.
Figure 12: iSCoBots learn to land in between the Flag to maximize its reward, when only $R_{1}^{exp}$ is provided.
In the main part of this manuscript, we address this problem on the game Skiing. To encourage the agent to go in between the flag, we add a reward that corresponds to the distance to both of the next flags (on the x axis):
$R^{\text{exp}}_{1}=D(Player,Flag1).x+D(Player,Flag2).x.$ (8)
This results in an agent learning to land close to the Flags (cf. Fig. 12)
To avoid this behavior, we add another signal, based on the speed of the agent:
$R^{\text{exp}}_{2}=V(Player).y.$ (9)
Misalignment: Reward shaping permits realigning the agent on the true task objectives, preventing undesirable behaviors resulting from misaligned optimization. Shortcut learning can be mitigated through the implementation of reward signals that necessitate the genuine objective’s consideration.

### v2

@@
 Abstract
-Reward sparsity, difficult credit assignment, and misalignment are only a few of the many issues that make it difficult, if not impossible, for deep reinforcement learning (RL) agents to learn optimal policies.
-Unfortunately, the black-box nature of deep networks impedes the inclusion of domain experts who could interpret the model and correct wrong behavior.
-To this end, we introduce Successive Concept Bottlenecks Agents (SCoBots), which make the whole decision pipeline transparent via the integration of consecutive concept bottleneck layers.
-SCoBots make use of not only relevant object properties but also of relational concepts.
-Our experimental results provide strong evidence that SCoBots allow domain experts to efficiently understand and regularize their behavior, resulting in potentially better human-aligned RL.
-In this way, SCoBots enabled us to identify a misalignment problem in the most simple and iconic video game, Pong, and resolve it.
+Goal misalignment, reward sparsity and difficult credit assignment are only a few of the many issues that make it difficult for deep reinforcement learning (RL) agents to learn optimal policies.
+Unfortunately, the black-box nature of deep neural networks impedes the inclusion of domain experts for inspecting the model and revising suboptimal policies.
+To this end, we introduce Successive Concept Bottleneck Agents (SCoBots), that integrate consecutive concept bottleneck (CB) layers.
+In contrast to current CB models, SCoBots do not just represent concepts as properties of individual objects, but also as relations between objects which is crucial for many RL tasks.
+Our experimental results [^1] Code available at https://github.com/k4ntz/SCoBots provide evidence of SCoBots’ competitive performances, but also of their potential for domain experts to understand and regularize their behavior.
+Among other things, SCoBots enabled us to identify a previously unknown misalignment problem in the iconic video game, Pong, and resolve it.
+Overall, SCoBots thus result in more human-aligned RL agents.
 Explaible AI (XAI), Reinforcement Learning, Concept Bottlenecks
 1 Introduction
-Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffering from a variety of issues that hinder them from learning optimal, generalisable action policies.
+Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalisable policies.
 Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023).
-A more recent, yet pressing issue is the goal misalignment problem (di Langosco et al., 2022).
-It occurs when an agent optimizes a different side-goal that is aligned with the target goal at training time, but not at test time (Koch et al., 2021; di Langosco et al., 2022).
-We discover that such misalignment can occur in the simplest and most iconic ALE (Mnih et al., 2013) game, Pong, in which an agent’s target goal is to play a ball game against an enemy.
-However, due to the nature of the automated enemy an agent can learn to focus on the position of the enemy paddle for anticipating where to position its own paddle for catching the ball, rather than the position of the ball itself.
-This can lead to failure of the agent when the enemy is occluded (cf.
+A more recent, yet pressing issue is the goal misalignment problem.
+It occurs when an agent optimizes a different side-goal, that is aligned with the original target goal during training (Koch et al., 2021), but not at test time.
+Such misalignments can be difficult to identify (di Langosco et al., 2022).
+For instance in this work we discover for the first time that such a misalignment can occur in the one of the oldest and most iconic Atari games, Pong (cf.
 Fig.
-1 (top)).
-This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020)) in which a model learns good decision rules for the provided training set that, however, fail to transfer to novel data sets.
-Importantly, if left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings.
+1, top).
+In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one (Mnih et al., 2013).
+Due to the automated enemy, that constantly follows the ball, agents can learn to focus on the position of the enemy paddle’s for placing their own paddle, rather than the position of the ball itself.
+This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020) in which a model learns decision rules for the provided training set, that fail to transfer to novel data sets.
+If left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings (Advani & Saxe, 2017; Zhang et al., 2021).
 Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision.
-Top: Deep RL agents trained on Pong produce high playing scores ( Refer to caption ) while importance map explanations suggest sensible underlying reasons of taking an action.
-However, when the enemy is hidden, the deep RL agent fails to even catch the ball ( Refer to caption ) without clear reasons.
+Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons of taking an action ( Refer to caption ).
+However, when the enemy is hidden, the deep RL agent fails to even catch the ball without clear reasons ( Refer to caption ).
 Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection , e.g., at a relational concept, but also action level.
@@
 In this way SCoBots can mitigate RL specific caveats like goal misalignment.
-A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI) (e.g.
-(Schramowski et al., 2020; Lapuschkin et al., 2019)) where the goal of XAI methods is to identify the reasons behind a model’s decisions (Guidotti et al., 2019; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023).
-Many methods of XAI however suffer from issues such as faithfulness (the explanation does not faithfully present a model’s underlying decision process)(Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020) and low-level semantics (explanations are made at the level of the input data e.g.
-pixels and not the relevant concepts that are extracted from these pixels)(Stammer et al., 2021; Kambhampati et al., 2021; Teso et al., 2023).
-A recent branch of research therefore focuses on developing models that provide inherent, concept-based explanations.
-Prominent examples of this idea are concept bottlenecks models (CBMs) which, among other things, have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020).
+A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI), whose aim is to identify the reasons behind a model’s decisions (Lapuschkin et al., 2019; Guidotti et al., 2019; Schramowski et al., 2020; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023).
+Many methods of XAI however suffer from issues such as a lack of faithfulness, i.e.
+the explanation does not faithfully present the model’s underlying decision process (Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020), and coarse, low-level explanation semantics.
+For example, importance-map explanations indicate the importance of an input element for a decision without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023).
+A recent branch of research therefore focuses on developing models that provide inherent concept-based explanations.
+Prominent examples of this idea are concept bottlenecks models (CBMs) which have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020).
 More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
-Despite the success and advantageous properties of CBMs up to now, no research has been performed on their integration into RL, particularly for tackling RL-specific issues discussed above.
-A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape etc… However, objects in RL environments such as ALE games interact with another (Kaiser et al., 2019).
-Therefore, in order to adequately solve such tasks additional relational concepts are necessary that, however, build upon these single object concepts.
-For this purpose, in this work we introduce Successive Concept Bottleneck Agents (SCoBots).
-SCoBots (cf.
+Despite the success and advantageous properties of CBMs for tackling RL-specific issues discussed above, no research has however been performed for their integration into RL.
+A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape, etc.
+However, objects in RL environments, even as simple as Atari games, interact with one another (Kaiser et al., 2019).
+In order to adequately solve such RL tasks via explicit concept representations, agents thus need to also extract necessary relational concepts.
+In this work we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf.
 Fig.
-1 (bottom)) hereby integrate multiple, successive concept bottlenecks into their decision processes where each bottleneck layer consists of concepts representations that integrate the concept representations of the previous bottleneck layer(s).
-Importantly, given a set of predefined concept functions that automatically set objects and concepts of initial bottleneck layers in relation to another the final bottleneck layer consists of valuable relational concept representations.
-SCOBots thus represent inherently explainable RL agents which in comparison to deep RL agents allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes, from single object concepts, to relational concepts to the actual action decision.
-In our evaluations on benchmark Atari games we provide experimental evidence not just of the performance abilities of SCoBots that is on par with deep RL agents.
-More importantly we showcase their ability to provide valuable explanations and their ability to mitigate a variety of RL specific issues, from reward sparsity to misalignment problems.
-Importantly, in the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents even on the arguably simplest of all Atari games, Pong.
-Utilizing the advantages of SCoBots this behaviour is, however, easily removed.
-Ultimately our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
+1, bottom) as an important extension of the original concept bottleneck approach to RL.
+SCoBots integrate successive concept bottlenecks into their decision processes where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer.
+Specifically, provided a set of predefined concept functions SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers.
+Finally, the set of relational and object concept representations are used for optimal action selection.
+SCoBots thus represent inherently explainable RL agents and in comparison to deep RL agents, they allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from the single object properties, through the relational concepts to the actual action decision.
+In our evaluations on the iconic Atari Learning Environments (Mnih et al., 2013) (ALE), we provide experimental evidence of the performance abilities of SCoBots that is on par with deep RL agents.
+More importantly, we showcase their ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts.
+In the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents, even on a simple game as Pong.
+By utilizing the interaction capabilities of SCoBots, this behaviour is easily removed.
+Ultimately, our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
 In summary our specific contributions are:
-We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their predictions.
-We show that SCoBots allow to follow their internal decision processes.
+We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their decision processes.
+We show that SCoBots allow to inspect their internal decision processes.
 We show that the inherent inspectable nature of SCoBots can be utilized for identifying previously unknown misalignment problems of RL agents.
-We show that they allow human interactions that lead to mitigating different RL specific learning issues.
+We show that they allow for human interactions for mitigating various RL specific learning issues.
 We proceed as follows.
 We first formally introduce SCoBots and discuss their specific properties.
-We continue with extensive experimental evaluations on Atari game environments via several analysis and metrics.
+We continue with experimental evaluations via several metrics and analysis.
 Before concluding, we touch upon related work.
@@
 2 Successive Concept Bottleneck Agents
-Let us start with providing the background notations on RL via markov decision processes (MDPs) as well as concept bottleneck models (CBMs).
+Let us start with providing the background notations of RL problems, modelled as markov decision processes (MDPs), and also of concept bottleneck models (CBMs).
 In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment.
-Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ with parameters $\theta$, are usually black-box models that process raw input states, e.g.
-a set of frames, and provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
-Concept bottleneck models (CBMs) initiate a learning process by extracting relevant concepts from raw input e.g.
-image data.
-These concepts can e.g.
-represent the color or position of a an Atari character.
-Subsequently, these extracted concepts are used for downstream tasks such as classification tasks.
-Formally, a bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathcal{X}$ into a concept representation c.
-Next, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output $y\in\mathcal{Y}$.
+Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parametrized by $\theta$, are usually black-box models that process raw input states, e.g.
+a set of frames, to provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
+Concept bottleneck models (CBMs) initiate a learning process by extracting relevant concepts from raw input (e.g.
+image data).
+These concepts can represent the color or position of a depicted object.
+Subsequently, these extracted concepts are used for downstream tasks such as image classification.
+Formally, a bottleneck model $g:x\rightarrow c$ transforms an input $x\in\mathbb{R}^{D}$ with $D$ dimensions into a concept representation $c\in\mathbb{R}^{k}$ (a vector of $k$ concepts).
+Next, the predictor network $f:c\rightarrow y$ uses this representation to generate the final target output (e.g.
+$y\in\mathbb{R}$ for classification problems).
 The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects.
-In contrast, RL agents must learn not just from static, labeled data but also through interaction with dynamic environments.
-RL tasks thus often necessitate relational reasoning, as they involve understanding the relationships between depicted instances that evolve through time and interact with one another.
+In contrast, RL agents must learn not just from static data but also through interaction with dynamic environments.
+RL tasks thus often require relational reasoning, as they involve understanding the relationships between instances that evolve through time and interact with another.
 This is crucial for learning effective policies in complex, dynamic environments.
-In this manuscript, we use specific fonts to distinguish object properties and relations.
+Note that, in the following descriptions, we use specific fonts to distinguish objects’ properties from their relations.
 2.1 Building inspectable ScoBots
 An underlying assumption of SCoBots is that their processing steps should be inspectable and understandable by a human user.
-This stands in contrast e.g.
-to unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020) where there is no guarantee for learning human-relatable concepts.
+This stands in contrast to e.g.
+unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020), where there is no guarantee for learning human-aligned concepts.
 SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations.
-These steps are described herafter and depicted in Fig.
+These steps are described hereafter and depicted in Fig.
 2.
-Similar to other RL agents, SCoBots process the last $n$ frames $\{x_{t}\}_{t=T-n}^{T}$, observed from the environment.
-For each frame, SCoBots employ an initial concept extraction method as is standard in other works on CBMs (Stammer et al., 2021; Koh et al., 2020) to extract objects and their properties.
+Similar to other RL agents, SCoBots process the last $n$ observed frames from a sequence of images, $s_{t}=\{x_{i}\}_{i=t-n}^{t}$.
+For each frame, SCoBots need an initial concept extraction method to extract objects and their properties, as in previous works on CBMs (Stammer et al., 2021; Koh et al., 2020).
 Specifically, we start from an initial bottleneck model, $\omega_{\theta_{1}}(\cdot)$, that is parameterized by parameter set, $\theta_{1}$.
-Given an image, $x_{t}$, this model provides a set of $c_{t}$ object representations, $\omega_{\theta_{1}}(x_{t})=\Omega_{t}=\{o_{t}^{j}\}_{j=1}^{c_{t}}$, where each object representation corresponds to a tensor of different extracted properties of that object, (e.g.
-the object category, position (i.e.
-its $x,y$ coordinates), size, … etc).
-As in previous works on CBMs this bottleneck model, $\omega_{\theta_{1}}(\cdot)$, can correspond to a model that was supervisedly pretrained for predicting these object properties from an image.
-One of the major differences to previous work on CBMs is that SCoBots now extract and utilize relational concepts that are based on the previously extracted objects and their properties.
-Formally, SCoBots utilize a successive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor.
-This model, $\mu$, is parameterized by a predetermined set of relational functions, $\mathcal{F}$, that are used to extract a set of $d_{t}$ relational concepts.
-These relations are based on each individual object or property for unary relations or combinations of these for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$.
-Without strong prior knowledge $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed.
-However, this set can easily be updated on the fly by a human user e.g.
-with additional, novel relation functions.
-Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (e.g.
-image $x_{t}$) and that is parameterized by $\theta_{2}$}.
-Up to now all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations.
-To guarantee understandability also within the action selection step of SCoBots we require to embed a specifically interpretable action selector.
-While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods they lack this required interpretability feature.
+Given a state, $s_{t}$, this model provides a set of $c_{i}$ object representations per frame, $\omega_{\theta_{1}}(s_{t})=\Omega_{t}=\{\{o_{i}^{j}\}_{j=1}^{c_{i}}\}_{i=t-n}^{t}$, where each object representation corresponds to a tensor of different extracted properties of that object, (e.g.
+its category, position (i.e.
+$x,y$ coordinates), etc.).
+As done in previous works on CBMs, the bottleneck model of our SCoBots, $\omega_{\theta_{1}}$, can correspond to a model that was supervisedly pretrained for extracting the objects and their properties from images.
+One of the major differences to previous work on CBMs is that SCoBots further extract and utilize relational concepts that are based on the previously extracted objects and their properties.
+Formally, SCoBots utilize a consecutive bottleneck model, $\mu_{\mathcal{F}}(\cdot)$, called the relation extractor.
+This model, $\mu$, is parameterized by a predetermined set of transparent relational functions, $\mathcal{F}$, used to extract a set of $d_{t}$ relational concepts.
+These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$.
+Without strong prior knowledge, $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed.
+However, this set can easily be updated on the fly by a human user with e.g.
+additional, novel relational functions.
+Note that $\mathcal{F}$ can include the identity, to let the relation extractor pass through initial object concepts from $\Omega_{t}$ to the relational concepts $\Gamma_{t}$.
+Finally, SCoBots employ an action selector, $\rho_{\theta_{2}}$, parameterized by $\theta_{2}$}, on the relational concepts to select the best action, $a_{t}$, given the initial state, $s_{t}$ (i.e.
+the set of frames).
+Up to now, all extracted concepts in SCoBots, both $\Omega_{t}$ and $\Gamma_{t}$, represent human-understandable concept representations.
+To guarantee understandability also within the action selection step of SCoBots, we need to embed an interpretable action selector.
+While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods, they lack this required interpretability feature.
 However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models.
-Thus, to trade-off the issues of flexibility and performance vs interpretability SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization.
+To trade-off these issues of flexibility and performance vs interpretability, SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization.
 After this, $\widetilde{\rho}_{\theta^{\prime}_{2}}$ is finally distilled into a decision tree $\rho_{\theta_{2}}$.
@@
 $s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}$ $\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$ (2)
-where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) are the policy parameters.
+where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters.
 For simplicity, we will discard the parameter notations in the rest of the manuscript.
-Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in App.A.6 and App.
-A.7.
+Further explanations of the input and output space of each module, as well as the properties and relational functions used in this work are provided in the appendix (cf.
+App.A.5 and App.
+A.6).
 Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently.
-This separation of training for different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem.
-In this way, the object-concept extractor, $\omega$, can be optimized by supervised (Redmon et al., 2016) or unsupervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods.
-On the other hand, the set of suitable relations, $\mathcal{F}$, as well which objects to apply them on, may be adapted via evolutionary algorithms (Peng et al., 2018; Yang et al., 2022), supervision from large language models (Wu et al., 2023), via prior task-knowledge (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b).
-Conclusively, the action selector, $\rho$ can be optimized based on the direct RL signal.
+Separating the training procedure of different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem.
 2.2 Guiding SCoBots
 The inspectable nature of SCoBots not only brings the benefit of improved human-understandability, but importantly allows for targeted human-machine interactions.
-In the following we describe two guidance approaches for interacting with the latent representations of SCoBots: concept pruning and object-centric rewarding.
+In the following, we describe two guidance approaches for interacting with the latent representations of SCoBots: concept pruning and object-centric rewarding.
+We refer to the revised SCoBot agents as guided SCoBots in the following.
 Concept pruning.
 The type and amount of concepts that are required for a successful policy may vary greatly across tasks.
-For instance, object colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman.
+For instance, the objects colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman.
 At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g.
@@
 Moreover, for a single RL task, the need for specific concepts might even change during training, e.g.
-in progressive environments where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2021).
-Specifically, SCoBot’s human-comprehensible concepts allow domain experts to prune unnecessary concepts.
-In this way expert users can guide the learning process towards relevant concepts.
+in progressive environments (where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024)).
+SCoBot’s human-comprehensible concepts therefore allow domain experts to prune unnecessary concepts.
+In this way, expert users can guide the learning process towards relevant concepts.
 Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$.
-Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, and (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$.
-iSCoBots thus formally refine Eq.
+Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, or (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$.
+Guided SCoBots thus formally refine Eq.
 2, leading to:
@@
 Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections.
-Human-normalized scores (red line) of deep and object-centric agents trained using PPO on 99 ALE environments.
-SCoBots obtain similar or better performance than the deep agents, showing object-centric agents can also learn to solve RL tasks while making use of human-understandable concepts (left).
-Guiding SCoBots allow to correct misalignment in Pong (center) and the ill-defined reward signal of Kangaroo (right).
+Human-normalized scores of deep and SCoBots (with and without guidance) agents trained using PPO on 99 ALE environments.
+SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left).
+Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level complition score of 100% on the intended goal’s evaluation in Kangaroo (right).
 To give brief examples of these four pruning possibilities we refer to Fig.
-2, focusing here on the orange subsections.
-Particularly, in Pong a user can remove objects (scores) or specific object properties (R,G,B values) which would result in a $\overline{\Omega_{t}}$.
-Second, the color relation, color($\cdot$), is not relevant for successfully playing Pong and can therefore be removed entirely from $\mathcal{F}$.
-Third, a user can state that the vertical distance function (dist($\cdot,\cdot$).y) should not be applied to the (player, enemy) input couple.
-These pruning actions result in the set of relational concepts $\overline{\Gamma_{t}}$.
+2, focusing here on the blue subparts.
+Particularly, in Pong a user can remove objects (e.g.
+scores) or specific object properties (e.g.
+R,G,B values) to obtain $\overline{\Omega_{t}}$.
+Second, the color relation, color($\cdot$), is irrelevant for successfully playing Pong and can therefore be removed from $\mathcal{F}$.
+Third, the vertical distance function (dist($\cdot,\cdot$).y) can be prevented from being applied to the (player, enemy) input couple.
+These pruning actions provide SCoBots with $\overline{\Gamma_{t}}$.
 Lastly, the only playable actions in Pong are UP and DOWN.
-To ease the search, the available FIRE, LEFT and RIGHT from the base Atari action space might be discarded as they are equivalent to NOOP (i.e.
-no action), resulting in $\overline{\mathcal{A}}$.
+To ease the policy search, the available FIRE, LEFT and RIGHT from the base Atari action space might be discarded to obtain $\overline{\mathcal{A}}$, as they are equivalent to NOOP (i.e.
+no action).
 Importantly, being able to prune concepts can help mitigate misalignement issues such as those of agents playing Pong (cf.
-Figure 1), where an agent can base it’s action selection on the enemy’s position.
-Specifically, by pruning the enemy position from the distance relation concept a user can guide the SCoBot to rely on relevant features for their decisions such as the ball’s position, rather than the spurious correlation with the enemy’s paddle.
+Fig.
+1), where an agent can base its action selection on the enemy’s position instead of the ball’s one.
+Specifically, pruning the enemy position from the distance relation concept enforces SCoBot agents to rely on relevant features for their decisions such as the ball’s position, rather than the spurious correlation with the enemy’s paddle.
 Object-centric feedback reward.
@@
 The object-centric and importantly concept-based approach of SCoBots allows to easily craft additional reward signals.
-Formally, one can use the extracted relational concepts to express new expert reward signal:
+Formally, one can use the extracted relational concepts to express a new expert reward signal:
 $R^{\text{exp}}(\Gamma_{t})$ $:=\sum_{g_{t}\in\Gamma_{t}}\alpha_{g_{t}}\cdot g_{t},$ (4)
-where $R^{\text{exp}}:\Gamma\longrightarrow\mathbb{R}$.
-Moreover, where $\Gamma_{t}=\mu(\omega(s_{t}))$ is the relational state, extracted by the relation extractor of a SCoBot.
+where $R^{\text{exp}}:\Gamma\longrightarrow\mathbb{R}$ and $\Gamma_{t}=\mu(\omega(s_{t}))$ is the relational state, extracted by the relation extractor.
 The coefficient $\alpha_{g_{t}}\in\mathbb{R}$ is used to penalize or reward the agent proportionally to relational concepts.
@@
 For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf.
-red components of Fig.
+“expert reward” arrow in Fig.
 2), with the intention of incentivizing the agent to maintain close proximity with the ball.
-As we show in our experimental evaluation, we can use this concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf.
+As we show in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf.
 App.
-A.9 for details on each problem).
-3 Experimental Evalautions
-In our evaluations, we investigate several properties and potential benefits of SCoBots. We specifically aim to answer the following research questions:
-  • Q1. Are SCoBot agents able to learn competitive policies on benchmark RL environments?
-  • Q2. Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes, e.g. policy misalignments?
-  • Q3. Can concept-based guidance let us mitigate RL caveats of SCoBot agents, e.g. policy misalignments?
-Figure 4: SCoBots allow to follow their decision process.
-Their object centricity and interpretable concepts allow us to follow their reasoning.
-The decision processes of SCoBots (extracted from the decision trees) are provided with their associated states for an guided SCoBot on Skiing (left) and a SCoBot on Pong (right).
-For examble, in this Skiing state, our iSCoBot selects RIGHT, as the distance between Player and the (left) Flag1 (on the $x$ axis) is high (bigger than $15$).
+A.8 for details on each problem).
+3 Experimental Evaluations
+In our evaluations, we investigate several properties and potential benefits of the SCoBot agents. We specifically aim to answer the following research questions:
+(Q1) Are concept based agents able to learn competitive policies on different RL environments?
+(Q2) Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes?
+(Q3) Can concept-based guidance help mitigate RL caveats of SCoBot agents, such as policy misalignments?
+Figure 4: SCoBots allow to follow their decision process, because of their object centricity and interpretable concepts.
+The decision processes of SCoBots (extracted from the decision trees) with the associated states from a guided SCoBot on Skiing (left), and from normal SCoBots on Pong (middle) and Kangaroo (right).
+For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$.
 This agent selects the correct action for the right reasons.
-Experimental setup: We evaluate SCoBots on $9$ Atari games (v5 version) from the ALE benchmark (Bellemare et al., 2012) (by far the most used RL framework (cf.
+Experimental setup: We evaluate SCoBots on $9$ Atari games (cf.
+Fig.
+3 (left) for complete list) from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf.
 App.
-A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game.
+A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game (i.e.
+still able to return the ball).
 For results based on achieved game scores we provide human normalized scores (following eq.
-5) that are averaged over $3$ seeds for each agent configuration.
-We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and specify these as neural agents in the following.
-Note that we evaluate these agents on the latest v5 version of the environments, leading to slightly worst results than (Schulman et al., 2017).
-We obtain similar results when evaluating on the old versions (cf.
+6) that are averaged over $3$ seeds for each agent configuration.
+We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following.
+Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in (Schulman et al., 2017), however we obtain similar results when evaluating on the old v4 versions (cf.
 Tab.
 1).
-For the SCoBots we assume access to a pretrained object extractor and thus provide our agents object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a).
-This allows us to focus our evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g.
+to prevent overfitting.
+All agents are trained for $20\text{M}$ frames under the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 implementation (Raffin et al., 2021) and its default hyperparameters (cf.
+Tab.
+2 in App.
+A.5).
+We focus our SCoBot evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g.
 (Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)).
-For all agents, we restrict training to $20\text{M}$ frames and use the Proximal Policy Optimization algorithm (PPO, (Schulman et al., 2017)), specifically the stable-baseline3 (Raffin et al., 2021) implementation and its default hyperparameters for both the deep neural and SCoBot agents (cf.
-Tab.
-2 in the Appendix).
-SCoBots’ inputs consist of vectors containing the values of concatenated concepts (i.e.
-properties and relations).
+We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a).
 Specifically, in our evaluations, the set of object properties consists of object class (e.g.
@@
 App.
-A.6.1 for implementation details).
-To distill SCoBots learned policy of SCoBots, we use the rule-based model extraction method from (Zarlenga et al., 2021).
+A.5.1 for implementation details).
+To distill SCoBots’ learned policies, we use a rule-based model extraction method, similar to (Zarlenga et al., 2021).
 For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf.
@@
 More details about the experimental setup and the hyperparameters are provided in App.
-A.6.
+A.5.
 SCoBots learn competitive policies (Q1).
-We present human-normalized (HN) scores of both SCoBot and neural agents trained on each investigated game individually in Fig.
+We present human-normalized (HN) scores of both SCoBot and deep agents trained on each investigated game individually in Fig.
 3 (left).
-Exact numerical results are provided in Table 1 (cf.
+Numerical values are provided in Tab.
+1 (cf.
 App.
-A.3).
-Overall, our results suggest that SCoBots provide competitive performances on every tested game despite the constraints of multiple bottleneck within their architectures.
-This shows algorithms where the object extraction is decoupled from the reasoning process that leads to the action selection can be developed.
-Other environments might require additional information extraction processes (e.g.
-MsPacman, in which the agent must navigate a maze), and we haven’t evaluated on such environments here.
-This constitute a limit to our SCoBots.
+A.7).
+We observe that SCoBot agents perform at least on par with deep agents on all games, even slightly outperform these on $5$ out of $9$ (namely, Asterix, Boxing, Kangaroo, Seaquest and Tennis).
+Our results suggest that SCoBots provide competitive performances on every tested game despite the constraints of multiple bottlenecks within their architectures.
+Overall, our experimental evaluations show that RL policies based on interpretable concepts extraction decoupled from the action selection process can, in principle, lead to competitive agents.
 Inspectable SCoBots’ to detect misalignments (Q2).
-A main target of developing ScoBots is to obtain competitive, yet transparent RL agents that allow for human users to identify an agent’s underlying reasoning process.
-Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allow to pinpoint not just the object properties, but importantly also the relational concepts and properties that are being used for the final action decision.
+The main target of developing ScoBots is to obtain competitive, yet transparent agents that allow for human users to identify their underlying reasoning process.
+Particularly, for a specific state, the inspectable bottleneck structure of SCoBots allows to pinpoint not just the object properties, but importantly the relational concepts being used for the final action decision.
 This is exemplified in Fig.
-4 for the games of Skiing and Pong (cf.
+4 on Skiing, Pong and Kangaroo (cf.
 App.
-A.5 for explanations of SCoBots on the remaining games).
-Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of the two games.
-For example, for the specific game state in Skiing, the SCoBot agent selects the action RIGHT as the best option, because the distance from its character to the left flag is larger than a specific value ($15$ pixels).
+A.4 for explanations of SCoBots on the remaining games).
+Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of each game.
+For example, for the specific game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels).
 Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
-A more striking display of the benefits the inherent explanations of SCoBot agents can be found in the Pong-based agent in Fig 4.
-Specifically, the provided explanation suggests that the agent is basing its decision on the vertical position of the enemy’s and of its own paddle (distance between the two paddles y direction).
-This, in turn, suggests that in fact the agent is ignoring the ball’s position, at least to a certain degree.
+A more striking display of the benefits of the inherent explanations of SCoBot agents is depicted by the Pong agent in Fig 4.
+Specifically, the provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis).
+In fact, this suggests that the agent is largely ignoring the ball’s position.
 Interestingly, upon closer inspection of the enemy movements, we observed that, previously unknown, the enemy is programmed to follow the ball’s vertical position (with a small delay).
-Thus, its vertical position is highly correlated with the enemy’s one (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf.
+Thus, the vertical positions of these two objects are highly correlated (Pearson’s and Spearman’s correlations coefficient above $99\%$, cf.
 App.
-A.8).
+A.7).
 Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.
-To validate these findings further we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (we refer to this environment as NoEnemy cf.
+To validate these findings further, we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (referred to as NoEnemy cf.
 Fig.
 1).
-We evaluate both the neural and SCoBot agents that were trained on the original Pong environment and observe catastrophic performance drops in HN scores (cf.
+We reevaluate the deep and SCoBot agents that were trained on the original Pong environment on NoEnemy and observe catastrophic performance drops in HN scores (cf.
 Fig.
-3 both of SCoBot and both neural PPO) at a level of random policies, when evaluated on this environment (NoEnemy).
-We have also run an DQN agent on this environments and observed the same results.
-The drop in neural agent performance is particularly striking as initial importance maps produced by these neural agents had highlighted all $3$ moving objects as relevant for their decisions (i.e.
-the player, ball, and enemy) (cf.
+3) at a level of random policies for both types of agents.
+Evaluations with different DQN agents lead to similar performance drops (cf.
+App.
+A.7).
+These drops are particularly striking as initial importance maps produced by PPO and DQN agents highlight all $3$ moving objects (i.e.
+the player, ball, and enemy) as relevant for their decisions (cf.
 Fig.1 top left and App.
-A.8, aligned with findings of (Weitkamp et al., 2018)).
-These importance maps suggest the agents had based their decision on right reasons.
-Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and but more importantly the granularity of such post-hoc explainability methods in the context of RL.
-It highlights the importance of inherently explainable RL models on the level of concepts as provided via SCoBots to identify potential issues such as goal misalignment.
+A.7) which is aligned with findings on importance maps of (Weitkamp et al., 2018).
+These maps suggest that the deep agents had based its decisions on right reasons.
+Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and granularity of such post-hoc explainability methods in the context of RL.
+Specifically, they highlight the importance of inherently explainable RL models on the level of concepts, as provided via SCoBots, to identify such potential issues.
 In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
 Guiding SCoBots to mitigate RL-specific caveats (Q3).
-We here make use of SCoBots interaction capabilities to address $4$ famous RL specific problems: goal misalignment which we had previously identified via SCoBot’s inherent, concept-based explanations and also sparse rewards, ill-defined rewards and difficult credit assignment.
-Realigning SCoBots: For mitigating identified SCoBots goal misalignment issues, we can correct the previously introduced misalignment, by simply removing (pruning) the enemy in the set of considered objects ($\overline{\Omega}$).
-Doing so, the enemy cannot be used for the action selection process.
-This leads to an agent that is able to play Pong and its NoEnemy version (guided SCoBot in Fig 3 (center)).
-This shows again that playing Pong without observing the enemy is achievable, and thus should also be the case for deep agents.
-Sparse Reward RL: Let us move on to the issue of sparse reward in the context of RL.
-In Pong, to score a point, the player needs to return the ball and have it go past the enemy.
-To do so, the player has to perform a spiky shot, obtained by touching the ball on one of its paddle’s sides (otherwise, the shot is flat and the enemy is easily catching it).
-The probability of such a shot is approximately $3.8\%$ (as shown in details in Appendix A.2).
-Thus, a randomly playing (or untrained) agent is only rewarded on average every $2200$ steps.
-Providing a SCoBot agent with an additional reward that is inversely proportional to the distance between its paddle and the ball incentivizes the agent to keep a vertical position close to the ball’s one, tripling the probability of a winning shot in our experiment (cf.
-App.
-A.2).
-Thus, the reward signal in Pong is low, but our example illustrates how the interpretable concepts allow to easily guide SCoBots, making up for the reward sparsity.
-Ill-defined reward: As previously explained, defining a reward upfront that incentivizes agents to learn the expected behavior can be difficult.
+We here make use of the interactive properties of SCoBots to address several famous RL specific problems: goal misalignment, ill-defined rewards, difficult credit assignment and reward sparsity.
+Realigning SCoBots: For mitigating goal misalignment issues of the SCoBot agents trained on Pong, we can simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$).
+By doing so, the enemy cannot be used for the action selection process.
+This leads to guided SCoBots that are able to play Pong and its NoEnemy version (cf.
+guided SCoBot in Fig 3 (center)).
+Furthermore, this shows that playing Pong without observing the enemy is achievable, by simply returning vertical shots, difficult for the enemy to catch.
+Ill-defined reward: Defining a reward upfront that incentivizes agents to learn an expected behavior is a difficult problem.
 An example of ill-defined reward (i.e.
 a reward that will lead to an unaligned behavior if the agent maximizes the return) is present in Kangaroo.
-According to the documentation of the game111www.retrogames.cz/play_195-Atari2600.php, “the mother kangaroo on the bottom floor tries to reach the top floor where her joey is being held captive by some monkeys”.
+According to the documentation of the game333www.retrogames.cz/play_195-Atari2600.php, “the mother kangaroo on the bottom floor tries to reach the top floor where her joey is being held captive by some monkeys”.
 However, punching the monkey enemies gives a higher cumulative reward than climbing up to save the baby.
-RL agents thus tend to learn to kick monkeys on the bottom floor and never go to the joey (cf.
+RL agents thus tend to learn to kick monkeys on the bottom floor rather than reaching the joey (cf.
 Fig.
 4).
-For revising such agents (i.e.
-incentivize them to reach the games’ original intended goal), we provide the SCoBot agents with an additional reward signal, based on the agent’s progression to the joey, (cf.
-App.
-A.9).
+For revising such agents, we provide an additional reward signal, based on the distance from the mother kangaroo to the joey (detailed in App.
+A.8).
 As can be seen in Fig.
-3 this reward allows us to obtain SCoBot agents that achieve high HN game scores Fig.
-3 (left) that indeed complete the level Fig.
-3 (right).
-In contrast the neural agents achieve high HN scores, but do not complete the levels, whereas unguided SCoBot agents achieve lower HN scores, but higher level completion ratios than the neural agents, but greatly lower ones than the guided SCoBot agents.
-Difficult Credit Assignment Problem: Lastly, the difficult credit assignment problem in RL is the challenge of correctly attributing credit or blame to past actions when the agent’s actions can have delayed and non-obvious effects on its overall performance.
-We illustrate this in the context of Skiing, where in standard configurations agents receive at each step a negative reward that corresponds to the number of seconds elapsed since the last steps (i.e.
+3 (right), this reward allows the guided SCoBots to achieve the originally intended goal by indeed completing the level.
+In contrast, deep agents and particularly unguided SCoBots achieve relatively high HN scores, but do not complete the levels.
+Difficult Credit Assignment Problem: The difficult credit assignment problem is the challenge of correctly attributing credit or blame to past actions when the agent’s actions have delayed or non-obvious effects on its overall performance.
+We illustrate this in the context of Skiing, where in standard configurations, agents receive at each step a negative reward that corresponds to the number of seconds elapsed since the last steps (i.e.
 varying between 3-3 and 7-7).
 This reward aims at punishing agents for going down the slope slowly.
-Additionally, the game keeps track of the number of flags that the agent has correctly traversed (displayed at the top of the screen, cf.
+Additionally, the game keeps track of the number of flag pairs that the agent has correctly traversed (displayed at the top of the screen, cf.
 Fig.
-4).
-A large positive reward, proportional to it, is attributed to the agent when it reaches the final state.
-Associating this large reward signal with the passed flags is extremely difficult, without prior knowledge on human skiing competitions.
-At each timestep, in the next evaluations we thus provide SCoBot agents with a reward that is proportional to the agent’s progression to the flags’ position to incentivize an agent to pass in between them as:
-$R^{\text{exp}}_{1}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}-D(Player,o)^{t-1}.$
-This resulted in an agent that learned to stop itself in between the flags (depicted in Fig.
-12), showing once again, how difficult it is to create a reward that would favor a specific behavior.
-To correct it, we therefore adjoined another signal to reward our SCoBot agents proportionally to the speed of the agent:
-$R^{\text{exp}}_{2}=V(Player).y.$
-These rewards are then summed to the environmental one.
-With this we obtain guided SCoBot agents that reach human scores, whereas the neural and non-interactive SCoBot agents perform worse than randomly (cf.
+4), and provide a large reward, proportional to this number, at the end of the episode.
+Associating this reward signal with the number of passed flags is extremely difficult, without prior knowledge on human skiing competitions.
+In the next evaluations, we provide SCoBots with another reward at each timestep, proportional to the agent’s progression to the flags’ position to incentivize the agent to pass in between them, as well as a signal rewarding for higher speed:
+$\begin{split}R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}&-D(Player,o)^{t-1}\\[-8.61108pt] &+V(Player).y.\end{split}$ (5)
+These rewards are further detailed in App.
+A.8.
+With this, we obtain guided SCoBot agents that reach human scores, whereas the deep and unguided SCoBot agents perform worse than random (cf.
 Fig.
 3 (left)).
-Overall, our experimental evaluations show that the not only are SCoBots on par with neural RL agents, but that importantly the concept-based and inspectable nature of SCoBots allows identify and revise several important RL specific caveats.
-In the course of our evaluations we could identify several previously unknown issues of RL agents on benchmark environments even as simple as Pong.
+Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf. App. A.8.3 for a detailled explanation).
+Overall, our experimental evaluations not only show that SCoBots are on par with deep agents in terms of performances, but that their concept-based and inspectable nature allows to identify and revise several important RL specific caveats.
+Importantly, in the course of our evaluations, we identified a previously unknown and critical misalignment of existing RL agents on the simple and iconic Pong environment, via the previously mentioned properties of SCoBots.
+4 Limitations
+In our evaluations, we made use of the integrated object extractor ($\omega$) of OCAtari which provides us with ground truth data.
+Such extractors can however also be optimized using supervised (Redmon et al., 2016) or self-supervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods.
+Furthermore, our set of relational functions, $\mathcal{F}$, is predefined and fixed.
+To eliminate the reliance on expert users, this set may alternatively be obtained using the supervision from large language models (Wu et al., 2023), via prior task-knowledge integration (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b).
+Other environments require additional information extraction processes.
+E.g.
+in MsPacman an agent must navigate a maze.
+Providing concept representations for such environments is an important step for future work.
+Finally, during our revision assessments, we presuppose that the user has adequate knowledge and positive intentions.
 5 Related Work
-The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as (Lampert et al., 2009) and (Kumar et al., 2009).
-A first systematic study of CBMs and their variations was delivered by (Koh et al., 2020) and (Stammer et al., 2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output.
+The basic idea of Concept Bottleneck Models (CBMs) can be found in work as early as Lampert et al.
+(2009) and Kumar et al.
+(2009).
+A first systematic study of CBMs and their variations was delivered by Koh et al.
+(2020) and Stammer et al.
+(2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output.
 Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$.
 Since the initial works on CBMs they have found utilization in several applications.
-(Antognini & Faltings, 2021) e.g.
+Antognini & Faltings (2021) e.g.
 apply CBMs to text sentiment classification.
-(Bahadori & Heckerman, 2020) propose a bias-removing method to tackle the problem of confounding concept correlations.
-(Chauhan et al., 2022) on the other hand query a domain expert for concept adaptation at prediction time.
-However, all of these works consider only single object concept representations and mainly focus on supervised learning settings.
+However, these works consider single object concept representations and focus on supervised learning.
 In fact, CBMs have found their way into RL only to a limited extent.
-(Zabounidis et al., 2023) and (Grupen et al., 2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance.
-(Zabounidis et al., 2023) further report better training stability and reduced sample complexity.
+Zabounidis et al.
+(2023) and Grupen et al.
+(2022) utilized CBMs in a multi-agent setting, where both identify improved interpretability through concept representations while maintaining competitive performance.
+Zabounidis et al.
+(2023) further report better training stability and reduced sample complexity.
 Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part.
 SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations.
-Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, (Guan et al., 2023) allow the user to design a reward function based on “behavioral attributes”, i.e.
-higher-level properties that occur not in a single state but over a period of time.
-Lastly, SCoBots not only separate state representation learning from policy search, as done by (Cuccu et al., 2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
+Similar to how SCoBots invite the user to reuse the high-level concepts to shape the reward signal, Guan et al.
+(2023) allow the user to design a reward function based on higher-level properties that occur over a period of time.
+Lastly, SCoBots not only separate state representation learning from policy search, as done by Cuccu et al.
+(2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
 Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022).
-For example, the true qualities of a good explanation are unclear and there is a lack of comparative explainability evaluation benchmarks.
-(Milani et al., 2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior.
+Milani et al.
+(2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior.
 Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods.
 According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining).
-Moreover, (Dazeley et al., 2021) suggest different “levels of explanation” based on the agent’s intentions.
-SCoBots especially facilitate the generation of zero-order explanations, i.e., reactive explanations that explain the agent’s reaction of an immediately perceived input.
-Overall, SCoBots transform the raw state $s$ into a symbolic state $s^{\prime}=\mu(\omega(s))$ in line with the proposal of Kambhampati et al..
-d’Avila Garcez et al.
-do an analogous symbolic extraction but resulting in a set of symbolic “sub-states” $s_{k}^{\prime}$ instead of $s^{\prime}$: Each $s_{k}^{\prime}$ represents the relative position of the agent to one interactable object.
-This way, however, solving one MDP results in solving a set of partially observable MDPs which is avoided in the case of SCoBots.
-The misalignment problem is related to shortcut learning, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023) to prototype-based models (Bontempelli et al., 2023) to reinforcement learning (RL) approaches (di Langosco et al., 2022).
+The misalignment problem is an RL instantiation of the shortcut learning problem, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023), prototype-based models (Bontempelli et al., 2023) and RL approaches (di Langosco et al., 2022).
 A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022).
-The authors (Nahian et al., 2021) align the agent to ethical values by introducing a second reward signal.
-In comparison, SCoBots aid to resolve RL specific learning issues such as the misalignment problem through inherent interpretability.
+Nahian et al.
+(2021) ethically align agents by introducing a second reward signal.
+In comparison, SCoBots aid to resolve RL specific issues such as the misalignment problem through inherent interpretability.
 6 Conclusion
-In this work, we have shown the benefits of concept-based models in RL tasks, specifically for identifying issues such as goal misalignment.
-Among other things our proposed Successive Concept Bottleneck agents specifically integrate automatically generated relational concept representations into their decision processes.
-With this we could identify previously unknown misalignment problems of deep RL agents in games as simple as Pong.
-Due to their inspectable and human-understandable nature of our proposed SCoBot agents these issues could not only be identified, but easily revised with minimal additional feedback.
+In this work, we have provided evidence for the benefits of concept-based models in RL tasks, specifically for identifying issues such as goal misalignment.
+Among other things our proposed Successive Concept Bottleneck agents integrate relational concepts into their decision processes.
+With this, we have exposed previously unknown misalignment problems of deep RL agents in a game as simple as Pong.
+SCoBot agents allowed us to revise this issue, as well as different RL specific caveats with minimal additional feedback.
 Our work thus represents an important step in developing aligned RL agents, i.e.
@@
 Achieving this is particularly valuable for applying RL agents in real-world settings where ethical and safety considerations are paramount.
-Avenues for future research are incorporating options, which represent higher-level actions as described in (Bacon et al., 2017), in order to make our (i)SCoBots more closely resemble human-like decision-making processes.
+Avenues for future research are incorporating a high level action bottleneck (Bacon et al., 2017).
 One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023).
-Additionally, we are considering the use of shallower decision trees, as proposed by (Broelemann & Kasneci, 2019), to enhance interpretability further in our RL agents.
-In our experiments we assume the initial object concept representations have been extracted.
-An interesting research question is further in how far the task reward signal can aid in learning such object-centric representations in the first place e.g.
-based on the MOC framework of Delfosse et al..
+Additionally, we are considering the use of shallower decision trees (Broelemann & Kasneci, 2019).
+An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023c) in the first place.
+Impact statement
+Our work aims at developing transparent RL agents, whose decision can be understood and revised to be aligned with the beliefs and values of a human user.
+We believe that such algorithms are critical to uncover and mitigate potential misalignments of AI systems.
+A malicious user can, however, utilize such approaches for aligning agents in a harmful way, thereby potentially leading to a negative impact on further users or society as a whole.
+Even so, the inspectable nature of transparent approaches will allow to identify such potentially harmful misuses, or hidden misalignment.
 Appendix A Appendix
-As mentioned in the main body, the appendix contains additional materials and supporting information for the following aspects: further information on the fact that Atari is the most common set of games (A.1), details on the reward sparsity in Pong (A.2), detailed numerical results (A.3), learning curves of our RL agents (A.4), the hyperparmeters used in this work (A.6), formal definitions on the SCoBots policies (A.7) and further details on the misalignment problem in Pong.
+As mentioned in the main body, the appendix contains additional materials and supporting information for the following aspects: further information on the fact that Atari is the most common set of games (A.1), details on the reward sparsity in Pong (A.8.3), detailed numerical results (A.2), learning curves of our RL agents (A.3), the hyperparmeters used in this work (A.5), formal definitions on the SCoBots policies (A.6) and further details on the misalignment problem in Pong.
 A.1 Atari games are most common set of environments
@@
 Figure 5: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from (Delfosse et al., 2023a).
-A.2 The reward sparsity of Pong.
-In the game Pong, the agent has a low chance of discovering positive reward.
-Its vertical position may vary between $34$ and $190$.
-Its paddle height is $16$, but it needs to shoot from the side of the paddle (the $3$ pixel of each border) to get a spiky shot that is likely to go beyond the opponent.
-If the balls arrives not close to the top or bottom borders, the enemy will still catch it, which has the probability of $\sim 22\%$ in our experiments.
-The probability of getting a successful shot is thus of $\sim 6/156*0.78=3\%$.
-The average number of corresponding tryouts for the agent to get rewarded is $\sum_{n=1}^{\infty}\left(n\times 0.03\times 0.97^{n-1}\right)=33.3$.
-The agent usually need $60$ steps (in average) from the initialization of the point to the reward attribution, hence, it will need $\sim 2200$ steps to be rewarded.
-This is consistent with the experiment depicted in Figure 6.
-In this figure, a random agent (original) needs in average $2230$ steps to observe reward.
-Figure 6: Expert user feedback allow for faster discovering of the reward signal in the sparse reward Pong environment.
-By providing the agent with a punishment proportional to the vertical distance between its paddle and the ball, we were able to obtain a winning shot every $\sim\!820$, multiplying by $2.7$ their occurrences.
-A.3 Detailed numerical results
+A.2 Detailed numerical results
 For completeness, we here provide the numerical results of the performances obtained by each agent type.
 For fair comparison, we reimplemeneted the deep PPO agents, and used the default hyperparameters for Atari from stable-baselines3 for the deep RL agents.
-A detailed overview of hyperparameters and their corresponding values can be found in A.6.
+A detailed overview of hyperparameters and their corresponding values can be found in A.5.
 All agents are trained on gymnasium’s atari learning environments using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank).
@@
 We use deterministic actions for both evaluation and testing stages.
-Game SCoBots SCoBots (guided) Neural(v5) Original reproduced (v4) Original Random Human
+Game
+SCoBots
+v5
+SCoBots
+guided-v5
+PPO
+ours-v5
+PPO
+ours-v4
+PPO
+Schulman et al.-v4
+Random Human
 Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433.3\mbox{\scriptsize$\pm 2772.9$}$ $4532$ $210$ $8503$
@@
 Table 1: ScoBots and guided ScoBots obtain similar results than deep agents averaged over 33 seeded runs.
-Deep refers to the agent that use a CNN instead of our ICB layers.
+Neural refers to the agent that use a CNN instead of our ICB layers.
 We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)).
@@
 […]
-A.4 Learning curves
+A.3 Learning curves
 Figure 6: Training Return over frames ($\times 10^{6}$) seen for   guided SCoBots,   SCoBots and   neural agents.
-A.5 Agent reasoning through the decision trees.
+A.4 Agent reasoning through the decision trees
 In this section, we provide more decision trees from SCoBots on states from Freeway and Bowling.
 Figure 7: Decision trees from Freeway and Bowling.
-A.6 Hyperparameters and experimental details
+A.5 Hyperparameters and experimental details
 Actors NN 8
@@
 Entropy coefficient $c_{2}$ $0.01$
-Table 2: PPO Hyperparameter Values. $\alpha$ linearly decreases from 11 to 0 over the course of training.
+Table 2: PPO Hyperparameter Values. $\alpha$ linearly decreases from $1$ to 0 over the course of training.
 All Experiments were run on a AMD Ryzen 7 processor, 64GB of RAM and one NVIDIA GeForce RTX 2080 Ti GPU.
@@
 (2018).
-The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation222github.com/openai/baselines.
+The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation444github.com/openai/baselines.
 This includes frame skipping of 4 and reward clipping.
@@
 A list of all hyperparameter values used is provided in Table 2.
-A.6.1 The properties and features used for SCoBots
+A.5.1 The properties and features used for SCoBots
 In this paper, we used different properties and functions (to create features).
@@
 velocity $v(x_{t},y_{t},x_{t-1},y_{t-1})$ velocity of object (vector)
-color $color(o_{1})$ the CSS2.1333https://www.w3.org/TR/CSS2/ color category (e.g. Red)
+color $color(o_{1})$ the CSS2.1555https://www.w3.org/TR/CSS2/ color category (e.g. Red)
 top k $Top_{k}((o_{1},...,o_{n}),k,class)$ only $k$ closest ($D_{e}$ to player) objects visible
@@
 Table 4: Feature selection and pruning for guided SCoBots. denotes the included features, whereas X the features that are pruned out.
-A.7 SCoBots policies: formal definitions
+A.6 SCoBots policies: formal definitions
 $\mathcal{S}$$\mathcal{O}\mathcal{F}\mathcal{R}$$\mathcal{A}$$\omega$$\mu$$\rho$
@@
 From this function perspective, the bottleneck consists of two stages: an object-centric bottleneck stage at $\mathcal{O}$ and a semantic bottleneck stage at $\mathcal{R}$.
-A.8 Pong misalignment problem.
-The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019).
-Here, an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time.
+A.7 Pong misalignment problem
+The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019), in which an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time.
 When the coin is repositioned at test time di Langosco et al.
-observed that trained deep RL agents avoid the coin and simply target the end of the level, indicating that the agents have in fact learned to follow a simpler side-goal rather than the underlying strategy.
-To find the correlation between the enemy’s and the ball’s y positions, we let an random agent play the game for $100000$ frames and collect the positions of the enemy and ball every $10$ frames.
+observed that trained deep RL agents avoid the coin and simply target the end of the level, indicating that agents in fact learn to follow a simpler side-goal rather than the underlying strategy.
+For the correlation between the enemy’s and the ball’s y positions, we let an random agent play the game for $100000$ frames and collect the positions of the enemy and ball every $10$ frames.
 We then compute different correlation coefficients, and obtain 99.6% as Pearson coefficient, 96.4% as Kendall coefficient and 99.5% as Spearman coefficient.
-Here are then importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2021).
+We also collected importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2024).
+Figure 9: More importance maps of DQN agents playing Pong, with ReLU (left) and rational activation functions (right).
 Figure 10: Importance maps of trained deep DQN agents playing Pong (left) show that all $3$ moving objects are important for the agent’s decision, suggesting that the agent takes aims at sending the ball past the enemy.
 The same trained agent completely outperform its enemy in the original environment while it cannot return the ball in a modified version of the game where the enemy is hidden.
-Figure 11: More importance maps of DQN agents playing Pong, with ReLU (left) and rational activation functions (right).
-A.9 RL specific caveats
+A.8 RL specific caveats
 In this section, we give more details on the different RL specific caveats, as well
-Sparse Reward: Some RL environments provide the agent with a rarely occurring reward signal, making it challenging to learn effectively.
-Such sparse reward environments are particularly present in the robotic field (Andrychowicz et al., 2017).
-Ill-defined Objectives: Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018).
+A.8.1 Ill-defined Objectives
+Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018).
 Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common (Andrychowicz et al., 2017).
@@
 16 reward = 0
-Difficult Credit Assignment: When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021).
+A.8.2 Difficult Credit Assignment
+When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021).
 This problem arises in games like chess, where reward is passed to the agent only when the game is over.
@@
 $R^{\text{exp}}_{1}=D(Player,Flag1).x+D(Player,Flag2).x.$ (9)
-This results in an agent learning to land close to the Flags (cf. Fig. 12)
-To avoid this behavior, we add another signal, based on the speed of the agent:
+This resulted in an agent that learned to stop itself in between the flags (depicted in Fig.
+11), showing once again, how difficult it is to create a reward that would favor a specific behavior.
+To correct it, we therefore adjoined another signal to reward our SCoBot agents proportionally to the speed of the agent:
 $R^{\text{exp}}_{2}=V(Player).y.$ (10)
-Misalignment: Reward shaping permits realigning the agent on the true task objectives, preventing undesirable behaviors resulting from misaligned optimization.
-Shortcut learning can be mitigated through the implementation of reward signals that necessitate the genuine objective’s consideration.
+A.8.3 The reward sparsity of Pong
+Let us move on to the issue of sparse reward in the context of RL.
+In Pong, to score a point, the player needs to return the ball and have it go past the enemy.
+To do so, the player has to perform a spiky shot, obtained by touching the ball on one of its paddle’s sides (otherwise, the shot is flat and the enemy is easily catching it).
+Its vertical position may vary between $34$ and $190$.
+Its paddle height is $16$, but it needs to shoot from the side of the paddle (the $3$ pixel of each border) to get a spiky shot that is likely to go beyond the opponent.
+If the balls arrives not close to the top or bottom borders, the enemy will still catch it, which has the probability of $\sim 22\%$ in our experiments.
+The probability of getting a successful shot is thus of $\sim 6/156*0.78=3\%$.
+The average number of corresponding tryouts for the agent to get rewarded is $\sum_{n=1}^{\infty}\left(n\times 0.03\times 0.97^{n-1}\right)=33.3$.
+The agent usually need $60$ steps (in average) from the initialization of the point to the reward attribution, hence, it will need $\sim 2200$ steps to be rewarded.
+This is consistent with the experiment depicted in Figure 12.
+In this figure, a random agent (original) needs in average $2230$ steps to observe reward.
+Figure 12: Expert user feedback allow for faster discovering of the reward signal in the sparse reward Pong environment.
+Providing a SCoBot agent with an additional reward that is inversely proportional to the distance between its paddle and the ball incentivizes the agent to keep a vertical position close to the ball’s one:
+$R^{\text{exp}}=D(Player,Ball).y$ (11)
+This extra reward signal lets a starting agent a winning shot every $\sim\!820$, multiplying by $2.7$ their occurrences.
+Thus, the reward signal in Pong is relatively sparse, but our example illustrates how the interpretable concepts allow to easily guide SCoBots, making up for the reward sparsity.
+Interestingly, this also giving a clearer incentive to the agent to concentrate on the ball’s position, and not on the enemy’s one.
+We observe that such agent do not rely on the enemy to master the game.
+Alternative techniques such as prioritized experience replay (Schaul et al., 2016) allow offline methods to learn in such environments, but providing a smoother reward signal is another elegant way to address sparsity.
+Note that sparser environments exists (such as robotics ones).
+A.8.4 Misalignment
+In our experimental evaluations, we used the concept based reward shaping to realign the agents on the true (target) task objectives, preventing undesirable behaviors resulting from misaligned optimization.
+Shortcut learning can be mitigated through the implementation of such reward signals that necessitate the genuine objective’s consideration.

### v4

@@
 Overall, SCoBots thus result in more human-aligned RL agents.
-Explaible AI (XAI), Reinforcement Learning, Concept Bottlenecks
 1 Introduction
-Reinforcement learning (RL) agents, particularly deep RL agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalisable policies.
-Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2023).
-A more recent, yet pressing issue is the goal misalignment problem.
-It occurs when an agent optimizes a different side-goal, that is aligned with the original target goal during training (Koch et al., 2021), but not at test time.
+Deep Reinforcement learning (RL) agents are prone to suffer from a variety of issues that hinder them from learning optimal or generalizable policies.
+Prominent examples are reward sparsity (Andrychowicz et al., 2017) and difficult credit assignment (Raposo et al., 2021; Wu et al., 2024).
+A more pressing issue is the goal misalignment problem.
+It occurs when an agent optimizes a different side-goal, aligned with the original target goal during training (Koch et al., 2021), but not at test time.
 Such misalignments can be difficult to identify (di Langosco et al., 2022).
-For instance in this work we discover for the first time that such a misalignment can occur in the one of the oldest and most iconic Atari games, Pong (cf.
+For instance, in this work we discover that such a misalignment can occur in the oldest and most iconic video game, Pong (cf.
 Fig.
-1, top).
-In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one (Mnih et al., 2013).
-Due to the automated enemy, that constantly follows the ball, agents can learn to focus on the position of the enemy paddle’s for placing their own paddle, rather than the position of the ball itself.
-This phenomenon is more generally described by the umbrella term shortcut learning (Geirhos et al., 2020) in which a model learns decision rules for the provided training set, that fail to transfer to novel data sets.
-If left unchecked, it may lead to a lack of model generalization and dangerous and unintuitive failures when such models are deployed in real-world settings (Advani & Saxe, 2017; Zhang et al., 2021).
+1).
+In Pong, the agent’s target goal is to catch a ball with its own paddle and to return it passed the enemy’s one.
+The enemy is programmed to constantly follow the ball, thus, agents can learn to focus on the position of the enemy paddle for placing their own, rather than the position of the ball itself.
+If left unchecked, such shortcut learning (Geirhos et al., 2020) may lead to a lack of model generalization and unintuitive failures e.g.
+at deployment time (Zhang et al., 2021).
 Figure 1: Successive Concept Bottlenecks Agents (SCoBots) allow for easy inspection and revision.
-Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons of taking an action ( Refer to caption ).
+Top: Deep RL agents trained on Pong produce high playing scores with importance map explanations that suggest sensible underlying reasons for taking an action ( Refer to caption ).
 However, when the enemy is hidden, the deep RL agent fails to even catch the ball without clear reasons ( Refer to caption ).
 Bottom: SCoBots, on the other hand, allow for multi-level inspection of the reasoning behind the action selection, e.g., at a relational concept, but also action level.
-Moreover, they allow for users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts.
-In this way SCoBots can mitigate RL specific caveats like goal misalignment.
-A recently popular approach for identifying such shortcut behaviour is via methods from the field of eXplainable AI (XAI), whose aim is to identify the reasons behind a model’s decisions (Lapuschkin et al., 2019; Guidotti et al., 2019; Schramowski et al., 2020; Ras et al., 2022; Roy et al., 2022; Saeed & Omlin, 2023).
-Many methods of XAI however suffer from issues such as a lack of faithfulness, i.e.
-the explanation does not faithfully present the model’s underlying decision process (Hooker et al., 2019; Chan et al., 2022; DeYoung et al., 2020), and coarse, low-level explanation semantics.
-For example, importance-map explanations indicate the importance of an input element for a decision without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023).
-A recent branch of research therefore focuses on developing models that provide inherent concept-based explanations.
-Prominent examples of this idea are concept bottlenecks models (CBMs) which have been shown to provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020).
-More importantly, CBMs allow to identify and effectively revise incorrect model behaviour on a higher, concept level (Stammer et al., 2021).
-Despite the success and advantageous properties of CBMs for tackling RL-specific issues discussed above, no research has however been performed for their integration into RL.
-A reason for this is that previous concept-based models depend only on concepts that are at a single object level, e.g., object color, shape, etc.
-However, objects in RL environments, even as simple as Atari games, interact with one another (Kaiser et al., 2019).
-In order to adequately solve such RL tasks via explicit concept representations, agents thus need to also extract necessary relational concepts.
-In this work we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf.
+Moreover, they allow users to easily intervene on them ( Refer to caption ) to prevent the agents from focusing on potentially misleading concepts.
+In this way, SCoBots can mitigate RL specific caveats like goal misalignment.
+Recently eXplainable AI (XAI) methods have emerged to detect such shortcut behavior, by identifying the reasons behind a model’s decisions (Schramowski et al., 2020; Roy et al., 2022; Saeed and Omlin, 2023).
+However, many XAI methods’ explanations do not faithfully present the model’s underlying decision process (Chan et al., 2022).
+For example, importance-map explanations indicate the importance of an input element without indicating why this element is important (Kambhampati et al., 2021; Stammer et al., 2021; Teso et al., 2023).
+A recent branch of research therefore focuses on models that provide inherent concept-based explanations.
+Prominent examples are concept bottlenecks models (CBMs), which provide predictive performances on par with standard deep learning approaches for supervised image classification (Koh et al., 2020).
+More importantly, CBMs allow to identify and revise incorrect model behavior on a concept level (Stammer et al., 2021).
+Contrary to exising CBMs’ fields of applications, RL requires relational reasoning (Kaiser et al., 2019).
+In this work, we therefore introduce Successive Concept Bottleneck Agents (SCoBots, cf.
 Fig.
-1, bottom) as an important extension of the original concept bottleneck approach to RL.
-SCoBots integrate successive concept bottlenecks into their decision processes where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer.
-Specifically, provided a set of predefined concept functions SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers.
+1, bottom) bringing the concept bottleneck approach to RL.
+SCoBots integrate successive concept bottlenecks into their decision processes, where each bottleneck layer provides concept representations that integrate the concept representations of the previous bottleneck layer.
+Specifically, provided a set of predefined concept functions, SCoBots automatically extract relational concept representations based on objects and their properties of the initial bottleneck layers.
 Finally, the set of relational and object concept representations are used for optimal action selection.
-SCoBots thus represent inherently explainable RL agents and in comparison to deep RL agents, they allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from the single object properties, through the relational concepts to the actual action decision.
-In our evaluations on the iconic Atari Learning Environments (Mnih et al., 2013) (ALE), we provide experimental evidence of the performance abilities of SCoBots that is on par with deep RL agents.
-More importantly, we showcase their ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts.
-In the course of our evaluations we identify previously unknown shortcut behaviour of deep RL agents, even on a simple game as Pong.
-By utilizing the interaction capabilities of SCoBots, this behaviour is easily removed.
-Ultimately, our work thereby illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues via relational concept based models.
-In summary our specific contributions are:
-We introduce Successive Concept Bottleneck agents (SCoBots) that integrate relational concept representations into their decision processes.
-We show that SCoBots allow to inspect their internal decision processes.
-We show that the inherent inspectable nature of SCoBots can be utilized for identifying previously unknown misalignment problems of RL agents.
-We show that they allow for human interactions for mitigating various RL specific learning issues.
-We proceed as follows.
-We first formally introduce SCoBots and discuss their specific properties.
-We continue with experimental evaluations via several metrics and analysis.
-Before concluding, we touch upon related work.
+SCoBots thus represent inherently explainable RL agents that, in comparison to deep RL agents, allow for inspecting and revising their learned decision policies at multiple levels of their reasoning processes: from single object properties, through relational concepts to the action selection.
+Our evaluations on the iconic Atari Learning Environments (ALE, Mnih et al.
+(2013)) provides experimental evidence that SCoBots perform on par with deep RL agents.
+More importantly, we showcase SCoBots’ ability to provide valuable explanations and the potential of mitigating a variety of RL specific issues, from reward sparsity to misalignment problems, via simple guidance from domain experts.
+We identify previously unknown shortcut behavior of deep agents, even on the simple Pong game.
+By utilizing the interaction capabilities of SCoBots, this behavior is easily corrected.
+Ultimately, our work illustrates the severity of goal misalignment issues in RL and the importance of being able to mitigate these and other RL specific issues, via relational concept based models.In summary, our contributions are:
+(i) We introduce Successive Concept Bottleneck agents (SCoBots).
+(ii) We show that SCoBots allow to inspect their internal decision processes.
+(iii) We show that their inherent inspectable nature can helps identifying unknown misalignments.
+(iv) We show that they allow for human interactions for mitigating various RL specific issues.
+We proceed as follows. We introduce SCoBots and discuss their specific properties. We continue with experimental evaluations and analysis. Before concluding, we touch upon related work.
 Figure 2: An overview of Successive Concept Bottlenecks Agents (SCoBots).
 SCoBots decompose the policy into consecutive interpretable concept bottlenecks (ICB).
-Objects and their properties are first extracted from the raw input, human-understandable functions are then employed to derive relational concepts and finally used to select an action.
+Objects and their properties are first extracted from the raw input, human-understandable functions are then employed to derive relational concepts, used to select an action.
 The understandable concepts enable interactivity.
@@
 2 Successive Concept Bottleneck Agents
-Let us start with providing the background notations of RL problems, modelled as markov decision processes (MDPs), and also of concept bottleneck models (CBMs).
-In this work, we represent an RL problem through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a}>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment.
-Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parametrized by $\theta$, are usually black-box models that process raw input states, e.g.
+In this work, we represent RL problems through the framework of a Markov Decision Process, $\mathcal{M}=<\mathcal{S},\mathcal{A},P_{s,a},R_{s,a},\gamma>$, with $\mathcal{S}$ as the state space, $\mathcal{A}$ the set of available actions, $P(s,a)$ the transition probability, and $R(s,a)$ the immediate reward function, obtained from the environment, and $\gamma$ the didscount factor.
+Classic deep RL policies, $\pi_{\theta}(s)=P(A=a|S=s)$ parameterized by $\theta$, are usually black-box models that process raw input states, e.g.
 a set of frames, to provide a distribution over the action space (Mnih et al., 2015; van Hasselt et al., 2016).
@@
 $y\in\mathbb{R}$ for classification problems).
-The main body of research on CBMs focuses on image classification tasks, where extracted concepts represent attributes of objects.
-In contrast, RL agents must learn not just from static data but also through interaction with dynamic environments.
+The main body of research on CBMs focuses on image classification tasks, where the extracted concepts represent attributes of objects.
+In contrast, RL agents must learn not just from static data, but also through interaction with dynamically evolving environments.
 RL tasks thus often require relational reasoning, as they involve understanding the relationships between instances that evolve through time and interact with another.
@@
 This stands in contrast to e.g.
-unsupervised CBM approaches (Jabri et al., 2019; Zhou et al., 2019; Srinivas et al., 2020), where there is no guarantee for learning human-aligned concepts.
+unsupervised CBM approaches Jabri et al.
+(2019); Zhou et al.
+(2019); Srinivas et al.
+(2020), where there is no guarantee for learning human-aligned concepts.
 SCoBots rather take a different approach by dividing the concept learning and RL-specific action selection into several inspectable steps that are grounded in human-understandable concept representations.
@@
 This model, $\mu$, is parameterized by a predetermined set of transparent relational functions, $\mathcal{F}$, used to extract a set of $d_{t}$ relational concepts.
-These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$.
+These relations are based on each individual object (and its properties) for unary relations or on combinations of objects for n-ary relations, and are denoted as $\mu_{\mathcal{F}}(\Omega_{t})=\Gamma_{t}=\{g_{t}^{k}\}_{k=1}^{d_{t}}$.
 Without strong prior knowledge, $\mathcal{F}$ can initially correspond to universal object relations such as distance and speed .
@@
 While neural networks, a standard choice in RL literature, are performative and easy to train using differentiable optimization methods, they lack this required interpretability feature.
-However, (Çağlar Aytekin, 2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models.
+However, Çağlar Aytekin (2022) have recently shown the equivalence between ReLU-based neural networks and decision trees, which in contrast represent inherently interpretable models.
 To trade-off these issues of flexibility and performance vs interpretability, SCoBots thus break down the action selection process by initially training a small ReLU-based neural network action selector, $\widetilde{\rho}_{\theta^{\prime}_{2}}$, via gradient-based RL optimization.
@@
 Lastly, note that one can add a residual link from the initial object concepts, $\Omega_{t}$, to the relational concepts, $\Gamma_{t}$ such that the action selector can, if necessary, also make decisions based on basic object properties, e.g., the height of an object.
-Overall, our approach preserves the MDP formulation used by classic deep approaches, but decomposes the classic deep policy (1) into a successive concept bottleneck one (2):
-$s_{t}$ $\xrightarrow{\pi_{\theta}}a_{t}$ (1)
-$s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}$ $\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$ (2)
-where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters.
+Overall, our approach preserves the MDP formulation used by classic deep approaches, but decomposes the classic deep policy $s_{t}\xrightarrow{\pi_{\theta}}a_{t}$ into a successive concept bottleneck one $s_{t}\xrightarrow{\omega_{\theta_{1}}}\Omega_{t}\xrightarrow{\mu_{\mathcal{F}}}\Gamma_{t}\xrightarrow{\rho_{\theta_{2}}}a_{t}$, where $\theta=(\theta_{1},\theta_{2},\mathcal{F}$) constitutes the set of policy parameters.
 For simplicity, we will discard the parameter notations in the rest of the manuscript.
@@
 A.6).
-Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable the optimization of each policy component independently.
-Separating the training procedure of different components (similar to that of (Koh et al., 2020)) particularly reduces the complexity of the overall optimization problem.
+Instead of jointly learning object detection, concept extraction, and policy search, SCoBots enable independent optimization of each policy component.
+Separating the training procedure of different components reduces the complexity of the overall optimization problem (Koh et al., 2020).
 2.2 Guiding SCoBots
@@
 Concept pruning.
-The type and amount of concepts that are required for a successful policy may vary greatly across tasks.
+The type and amount of concepts that are required for a successful policy may vary across tasks.
 For instance, the objects colors are irrelevant in Pong but required to distinguish vulnerable ghosts from dangerous ones in MsPacman.
-At the same time overloading the action selector with too many irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g.
+However, overloading the action selector with irrelevant concepts, whether these are object properties ($\Omega_{t}$) or relational concepts ($\Gamma_{t}$), can lead to difficult optimization (e.g.
 the agent focusing on noise in unimportant states) as well as difficult inspection of the decision tree ($\rho$).
-Moreover, for a single RL task, the need for specific concepts might even change during training, e.g.
-in progressive environments (where agents are required to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024)).
-SCoBot’s human-comprehensible concepts therefore allow domain experts to prune unnecessary concepts.
-In this way, expert users can guide the learning process towards relevant concepts.
-Formally, users can (i) select a subset of the object property concepts, $\overline{\Omega}$.
-Additionally, by (ii) selecting a subset of the relational functions, $\overline{\mathcal{F}}$, or (iii) specifying which objects specific functions should be applied on, experts can implicitly define the relational concepts subset, $\overline{\Gamma}$.
-Guided SCoBots thus formally refine Eq.
-2, leading to:
+Moreover, for a single RL task, the need for specific concepts might even change during training, in e.g.
+progressive environments (where agents need to master early stages before being provided with additional, more complex tasks (Delfosse et al., 2024c)).
 […]
-Furthermore, users can (iv) prune out redundant actions resulting in a new action space $\overline{\mathcal{A}}$.
-Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections.
-Human-normalized scores of deep and SCoBots (with and without guidance) agents trained using PPO on 99 ALE environments.
-SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left).
-Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level complition score of 100% on the intended goal’s evaluation in Kangaroo (right).
 To give brief examples of these four pruning possibilities we refer to Fig.
@@
 In practice, this expert reward signal can lead to guiding the agent towards focusing on relevant concepts, but also help smoothing sparse reward signals, which we will discuss further in our evaluations.
-For example, experts have the ability to impose penalties, based on the distance between the agent and objects (cf.
+For example, one can impose penalties, based on the distance between the agent and objects (cf.
 “expert reward” arrow in Fig.
 2), with the intention of incentivizing the agent to maintain close proximity with the ball.
-As we show in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf.
+As shown in our experimental evaluation, we can use this concept pruning and concept based reward shaping to easily address many RL specific caveats such as reward sparsity, ill-defined objectives, difficult credit assignment, and misalignment (cf.
 App.
@@
 3 Experimental Evaluations
-In our evaluations, we investigate several properties and potential benefits of the SCoBot agents. We specifically aim to answer the following research questions:
+In our evaluations, we investigate several properties and potential benefits of the transparent SCoBot agents. We specifically aim to answer the following research questions:
 (Q1) Are concept based agents able to learn competitive policies on different RL environments?
-(Q2) Does the inherent inspectable nature of SCoBots allow to detect potential issues in their decision processes?
-(Q3) Can concept-based guidance help mitigate RL caveats of SCoBot agents, such as policy misalignments?
-Figure 4: SCoBots allow to follow their decision process, because of their object centricity and interpretable concepts.
-The decision processes of SCoBots (extracted from the decision trees) with the associated states from a guided SCoBot on Skiing (left), and from normal SCoBots on Pong (middle) and Kangaroo (right).
-For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$.
-This agent selects the correct action for the right reasons.
+(Q2) Does the inspectable nature of SCoBots allow to detect issues in their decision processes?
+(Q3) Can concept-based guidance help mitigate common RL caveats, such as policy misalignments?
+(Q4) Can SCoBots learn with imperfect object extraction methods?
+(Q5) How crucial is the the relation extractor for SCoBots performances?
 Experimental setup: We evaluate SCoBots on $9$ Atari games (cf.
 Fig.
-3 (left) for complete list) from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf.
+3 from the Atari Learning Environments (Bellemare et al., 2012) (by far the most used RL framework (cf.
 App.
-A.1)), as well as a variation of Pong where the enemy is not visible yet active in the game (i.e.
-still able to return the ball).
-For results based on achieved game scores we provide human normalized scores (following eq.
-6) that are averaged over $3$ seeds for each agent configuration.
-We compare our SCoBot agents to deep agents with the classic convolutional network introduced by (Mnih et al., 2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following.
-Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in (Schulman et al., 2017), however we obtain similar results when evaluating on the old v4 versions (cf.
+A.1), as well as the HackAtari modified (Delfosse et al., 2024a) Pong environments, where the enemy is not visible yet active (NoEnemy), and where the enemy stops moving after returning the ball (LazyEnemy).
+We provide human normalized scores (following eq.
+3) that are averaged over $3$ seeds for each agent configuration.
+We compare our SCoBot agents to deep agents with the classic convolutional network introduced by Mnih et al.
+(2015), that process a stack of $4$ black and white frames and denote these as deep agents in the following.
+Note, that we evaluate all agents on the latest v5 version of the environments [^2] This leads to deep agents overall slightly worse than in Schulman et al.
+(2017), however we obtain similar results when evaluating on the old v4 versions (cf.
 Tab.
@@
 We focus our SCoBot evaluations on the more interesting aspects of the agent’s reasoning process underlying the action selection, rather than the basic object identification step (which has been thoroughly investigated in previous works e.g.
-(Koh et al., 2020; Stammer et al., 2021; Wu et al., 2023)).
-We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2023a).
+(Koh et al., 2020; Stammer et al., 2021; Wu et al., 2024)).
+We thus assume access to a pretrained object extractor and provide our agents with object-centric descriptions of these states ($\Omega_{t}$) based on the information from OCAtari (Delfosse et al., 2024b).
 Specifically, in our evaluations, the set of object properties consists of object class (e.g.
@@
 A.5.1 for implementation details).
-To distill SCoBots’ learned policies, we use a rule-based model extraction method, similar to (Zarlenga et al., 2021).
+To distill SCoBots’ learned policies, we use the decision-tree extraction algorithm VIPER (Bastani et al., 2018).
 For the human-guided SCoBot evaluations (denoted as (guided) SCoBots in our evaluations), we illustrate human guidance and prune out the concepts that we consider unimportant to master the game (cf.
@@
 Furthermore, to mitigate RL specific problems, we also provide SCoBot agents with simulated human-expert feedback signals, the details of which we describe at the relevant sections below.
-More details about the experimental setup and the hyperparameters are provided in App.
+More details about the setup and the hyperparameters are in App.
 A.5.
+In our evaluation, SCoBots’ training is slightly faster than deep agents’ one (cf.
+Appendix A.9).
+Figure 3: Object-centric agents can master different Atari environments and interactive SCoBots allow for corrections.
+Human-normalized scores of different agents trained using PPO on 99 ALE environments, including deep agents (i.e.
+using CNNs), guided decision tree policy (SCoBots), their neural object-centric baseline (NN-), and these baselines without guidance (NG).
+SCoBots obtain similar or better scores than the deep agents, showing that object-centric agents can also solve RL tasks while making use of human-understandable concepts (left).
+Guiding SCoBots allow to correct misalignment in Pong (center) and to obtain the originally intended agents, depicted by a level completion score of 100% on the intended goal’s evaluation in Kangaroo (right).
 SCoBots learn competitive policies (Q1).
@@
 Overall, our experimental evaluations show that RL policies based on interpretable concepts extraction decoupled from the action selection process can, in principle, lead to competitive agents.
+Figure 4: Interpretable SCoBots allow to follow their decision process, thanks to their interpretable concepts.
+The states and associated decision processes of SCoBots (extracted from the decision trees) on Skiing (left), and from unguided SCoBots on Pong (middle) and Kangaroo (right).
+For example, in this Skiing state, our SCoBot selects RIGHT, as the signed distance between Player and the (left) Flag1 (on the $x$ axis) is bigger than $15$.
+This agent selects the correct action for the right reasons.
 Inspectable SCoBots’ to detect misalignments (Q2).
@@
 Here we highlight the decision-making path (from SCoBots’ decision tree based policies), at specific states of each game.
-For example, for the specific game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels).
+For example, for the game state extracted from Skiing, the SCoBot agent selects RIGHT as the best action, because the signed distance from its character to the left flag is larger than a specific value (+15+15 pixels).
 Given the nature of the game this inherent explanation suggests that the agent is indeed basing its selection process on relevant relational concepts.
 A more striking display of the benefits of the inherent explanations of SCoBot agents is depicted by the Pong agent in Fig 4.
-Specifically, the provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis).
+The provided explanation suggests that the agent is basing its decision on the vertical positions of the enemy and of its own paddle (distance between the two paddles on the $y$ axis).
 In fact, this suggests that the agent is largely ignoring the ball’s position.
@@
 A.7).
-Moreover, the ball’s rendering can lead to visual flickering, explaining why SCoBots base their decision on the potentially more reliable feature of the enemy paddle’s vertical position.
-To validate these findings further, we perform evaluations on a modified Pong environment in which the enemy is invisible, yet playing (referred to as NoEnemy cf.
+Moreover, the ball’s rendering contains flickering, explaining why SCoBots base their decision on this potentially more reliable feature, the enemy paddle’s vertical position.
+To validate these findings further, we perform evaluations on $2$ modified Pong environments in which (i) the enemy is invisible, yet playing (NoEnemy, cf.
 Fig.
-1).
+1), and on one environment where the enemy is not moving after returning the ball (LazyEnemy).
 We reevaluate the deep and SCoBot agents that were trained on the original Pong environment on NoEnemy and observe catastrophic performance drops in HN scores (cf.
@@
 Fig.1 top left and App.
-A.7) which is aligned with findings on importance maps of (Weitkamp et al., 2018).
+A.7), aligned with findings on importance maps of Weitkamp et al.
+(2018).
 These maps suggest that the deep agents had based its decisions on right reasons.
-Our novel results on the NoEnemy environment however greatly calls to question the faithfulness and granularity of such post-hoc explainability methods in the context of RL.
-Specifically, they highlight the importance of inherently explainable RL models on the level of concepts, as provided via SCoBots, to identify such potential issues.
-In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottleneck structure.
+Our novel results on the NoEnemy and LazyEnemy environments, however, greatly calls to question the faithfulness and granularity of such post-hoc explainability methods.
+They highlight the importance of inherently transparent RL models on the level of concepts, as provided via SCoBots, to identify such potential issues.
+In the following, we will investigate how to mitigate the discovered issues via SCoBot’s bottlenecks.
 Guiding SCoBots to mitigate RL-specific caveats (Q3).
 We here make use of the interactive properties of SCoBots to address several famous RL specific problems: goal misalignment, ill-defined rewards, difficult credit assignment and reward sparsity.
-Realigning SCoBots: For mitigating goal misalignment issues of the SCoBot agents trained on Pong, we can simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$).
-By doing so, the enemy cannot be used for the action selection process.
-This leads to guided SCoBots that are able to play Pong and its NoEnemy version (cf.
-guided SCoBot in Fig 3 (center)).
+Realigning SCoBots: To mitigate the goal misalignment issues of the SCoBots trained on Pong, we simply remove (prune) the enemy from the set of considered objects ($\overline{\Omega}$).
+Thus, the enemy cannot be used for the action selection process.
+This leads to SCoBots that are able to play Pong and its NoEnemy version (cf.
+SCoBot in Fig 3 (center)).
 Furthermore, this shows that playing Pong without observing the enemy is achievable, by simply returning vertical shots, difficult for the enemy to catch.
@@
 In the next evaluations, we provide SCoBots with another reward at each timestep, proportional to the agent’s progression to the flags’ position to incentivize the agent to pass in between them, as well as a signal rewarding for higher speed:
-$\begin{split}R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}&-D(Player,o)^{t-1}\\[-8.61108pt] &+V(Player).y.\end{split}$ (5)
+$R^{\text{exp}}=\smashoperator[]{\sum_{o\in\{Flag_{1},Flag_{2}\}}^{}}D(Player,o)^{t}-D(Player,o)^{t-1}+V(Player).y.$ (2)
 These rewards are further detailed in App.
 A.8.
-With this, we obtain guided SCoBot agents that reach human scores, whereas the deep and unguided SCoBot agents perform worse than random (cf.
+They allow guided SCoBots to reach human scores, whereas the deep and unguided SCoBots perform worse than random (cf.
 Fig.
 3 (left)).
-Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf. App. A.8.3 for a detailled explanation).
+Note that providing additional reward signals, as done above, obviously also allows to mitigate reward sparsity, as we illustrate on Pong (cf.
+App.
+A.8.3 for a detailed explanation).
+Figure 5: SCoBots can learn with noisy object detectors, transparent SCoBots rely on relations.
+Final human normalized scores (with stds) comparing SCoBots and the object-centric neural baselines (NN-SCoBots), with and without relations.
+We also provide the scores of NN-SCoBots that learned on noisy environments.
+The noise only noticeably affects the agents on Kangaroo.
+Ablating the relations is harmless on NN-SCobots, as neural networks can recompute them, but impacts SCoBots performances on 66 games.
+(For better visualization, we used a human score of $100$ in Boxing.)
+Object-centric agents can learn with imperfect object extractors (Q4).
+To test object-centric agents’ ability to work in more realistic environments, we have tested if they can learn viable policies with imperfect object extraction methods.
+Based on Delfosse et al.
+(2023b)’s detection results, we added a $5\%$ misdetection probability and a Gaussian noise (of 0 mean and $3$ pixels of standard deviation on each axis).
+As depicted in Figure 5, NN-SCoBots (i.e.
+neural network based object-centric agents, using relations) learn comparable policies on all games but Kangaroo, demonstrating their ability to mitigate suboptimal object extraction methods.
+Implementing robustness techniques, such as Kalman filters, would help to further push the performances of such agents.
+Transparent SCoBots benefit from explicit relations (Q5).
+While the ablation of the relation extractor only has a significant impact on the neural-based object-centric agents on $1$ out of $9$ games (Seaquest), it deteriorates the decision tree based SCoBots on $6$ environments.
+This is due to the fact that relations can be implicitly recomputed within a neural network, but not within decision trees.
+As shown by Kohler et al.
+(2024), the use of the distance relation (on a specific axis) allows for performing agents with compact decision tree-based policies.
+Furthermore, even if some relations can be rediscovered and implicitly encoded within the decision trees, the lack of explicit relational representations can reduce the interpretability of the agents, and will prevent the experts from using these within their guidance.
 Overall, our experimental evaluations not only show that SCoBots are on par with deep agents in terms of performances, but that their concept-based and inspectable nature allows to identify and revise several important RL specific caveats.
@@
 4 Limitations
-In our evaluations, we made use of the integrated object extractor ($\omega$) of OCAtari which provides us with ground truth data.
-Such extractors can however also be optimized using supervised (Redmon et al., 2016) or self-supervised (Lin et al., 2020; Delfosse et al., 2023c) object detection methods.
-Furthermore, our set of relational functions, $\mathcal{F}$, is predefined and fixed.
-To eliminate the reliance on expert users, this set may alternatively be obtained using the supervision from large language models (Wu et al., 2023), via prior task-knowledge integration (Reid et al., 2022), or neural guidance from a pretrained fully deep agent (Delfosse et al., 2023b).
+On the use of OCAtari.
+To limit our resource consumption, we made use of the quasi-perfect object extractor ($\omega$) of OCAtari, which efficiently extract objects from the RAM.
+We added an ablation to simulate potential imperfect detection capabilities of Atari object extraction methods.
+Such extractors can be optimized using supervised (Redmon et al., 2016; Locatello et al., 2020) or self-supervised (Lin et al., 2020; Delfosse et al., 2023b) object detection methods.
+This last work showcase that unsupervised object extraction methods can replace OCAtari at test time, however leading to performance drops.
+Grandien et al.
+(2024) have further showcase training RL agents with such pretrained object extractors further improve the object-centric RL agents performances.
+The limit of object-centricity.
 Other environments require additional information extraction processes.
 E.g.
-in MsPacman an agent must navigate a maze.
-Providing concept representations for such environments is an important step for future work.
-Finally, during our revision assessments, we presuppose that the user has adequate knowledge and positive intentions.
+in MsPacman, an agent must navigate a maze.
+Extending the concept representations to cover such concepts in maze or platform environments is an important step for future work.
+Other representations could allow for the integration of e.g.
+path finding methods such as the A* algorithm.
+Training time of SCoBots.
+Thanks to the efficient OCAtari object extraction, SCoBots required in average $7.5$ hours of training time, while deep agents needed $10.8$ hours (cf.
+Appendix A.9).
+All SCoBots variations require less training time compared to deep agents on environments with few objects (e.g.
+Boxing, Pong, Tennis).
+Currently, at every step, the concept bottleneck values are calculated sequentially in a single process on the CPU, leaving significant room for training time improvement by optimizing these computations (bringing them to GPUs).
+Thus, in environments with many objects, e.g.
+Kangaroo, the training time of unguided SCoBots exceeded that of its deep agent counterpart.
+Attention on relevant objects could be used to further save computational resources.
 5 Related Work
@@
 (2009).
-A first systematic study of CBMs and their variations was delivered by Koh et al.
-(2020) and Stammer et al.
-(2021) which described a CBM as a two-stage model that first computes intermediate concept representations before generating the final task output.
-Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Stammer et al., 2022; Steinmann et al., 2023; Sawada & Nakamura, 2022; Marconato et al., 2022; Lage & Doshi-Velez, 2020), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$.
+A first systematic study of CBMs was delivered by Koh et al.
+(2020), followed by Stammer et al.
+(2021), who described CBMs as two-stage models that first computes intermediate representations used for the final task output.
+Where learning valuable initial object concept representations without strong supervision is still a tricky and open issue for concept-based models (Lage and Doshi-Velez, 2020; Stammer et al., 2022; Sawada and Nakamura, 2022; Marconato et al., 2022; Steinmann et al., 2023; Stammer et al., 2024), receiving relational concept representations in SCoBots is performed automatically via the function set $\mathcal{F}$.
 Since the initial works on CBMs they have found utilization in several applications.
-Antognini & Faltings (2021) e.g.
-apply CBMs to text sentiment classification.
+Antognini and Faltings (2021) e.g.
+apply CBMs to text sentiment classification, and Kraus et al.
+(2024) to time series’ analysis.
 However, these works consider single object concept representations and focus on supervised learning.
+The ability of users to revise concepts and decisions has also been parallelly shown by Friedrich et al.
+(2023).
 In fact, CBMs have found their way into RL only to a limited extent.
@@
 (2023) further report better training stability and reduced sample complexity.
-Their extension includes an optional “residual” layer which passes additional, latent information to the action selector part.
+Their extension includes an optional “residual" layer which passes additional, latent information to the action selector part.
 SCoBots omit such a residual component for the sake of human-understandability, yet offer the flexibility to the user to modify the concept layer via updating $\mathcal{F}$ for improving the model’s representations.
@@
 (2020), but also enforce object-centricity, putting interpretability requirements on the consecutive feature spaces.
-Explainable RL (XRL) is an extensively surveyed XAI subfield (Milani et al., 2023; Dazeley et al., 2022; Krajna et al., 2022) with a wide range of yet unsolved issues (Vouros, 2022).
+Explainable RL (XRL) is an extensively surveyed XAI field (Dazeley et al., 2022; Krajna et al., 2022; Milani et al., 2023) with a wide range of unsolved issues (Vouros, 2022).
 Milani et al.
-(2023) introduce a three-category taxonomy for XRL methods: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present instances of the past experience or components of the MDP that imply the current behavior, and (3) policy-level methods which describe the agent’s long-term behavior.
-Based on this, SCoBots extract relations from low-level features making high-level information available to explanations and thereby support feature importance methods.
-According to (Qing et al., 2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining).
+(2023) introduce a taxonomy for XRL methods, with: (1) feature importance methods that generate explanations that point out decision-relevant input features, (2) learning process & MDP methods which present which past experience or MDP components affect the policy, and (3) policy-level methods, describing the agent’s long-term behavior.
+Based on this, SCoBots extract relations from low-level features, making high-level information available to explanations and thereby support feature importance methods.
+According to Qing et al.
+(2022)’s categorization of XRL frameworks, SCoBots are “model-explaining” (in contrast to reward-, state-, and task-explaining).
+Other XRL methods rely on LLM to explain the policy (Luo et al., 2024) decision trees (Fuhrer et al., 2024; Marton et al., 2024), logic (Jiang and Luo, 2019; Kimura et al., 2021; Delfosse et al., 2023a; Sha et al., 2024) or programs (Verma et al., 2018; Trivedi et al., 2021; Cao et al., 2022; Kohler et al., 2024; Wüst et al., 2024) to encode transparent policies.To overcome the potentially unavailable concepts necessary to learn symbolic policies, Shindo et al.
+(2024) learn a mixture of symbolic and neural policies.
+Finally, concepts have further been used to derive reward from context using LLMs, as we did for Kangaroo and Skiing (Kwon et al., 2023; Kaufmann et al., 2024; Wu, 2024; Shen et al., 2024).
 The misalignment problem is an RL instantiation of the shortcut learning problem, a frequently studied failure mode that has been identified in models and datasets across the spectrum of AI from deep networks (Lapuschkin et al., 2019; Schramowski et al., 2020) to neuro-symbolic (Stammer et al., 2021; Marconato et al., 2023), prototype-based models (Bontempelli et al., 2023) and RL approaches (di Langosco et al., 2022).
-A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold & Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022).
+A misaligned RL agent, first empirically studied by (Koch et al., 2021) represents a serious issue, especially when it is misaligned to recognized ethical values (Arnold and Kasenberg, 2017) or if the agent has broad capabilities (Ngo, 2022).
 Nahian et al.
@@
 Avenues for future research are incorporating a high level action bottleneck (Bacon et al., 2017).
-One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by (Wu et al., 2023).
-Additionally, we are considering the use of shallower decision trees (Broelemann & Kasneci, 2019).
-An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023c) in the first place.
+One can also incorporate attention mechanisms into RL agents, as discussed in (Itaya et al., 2021), or use language models (and e.g., the game manuals) to generate the additional reward signal, as done by Wu et al.
+(2024).
+Additionally, we are considering the use of shallower decision trees (Broelemann and Kasneci, 2019).
+An interesting research question is how far the task reward signal can aid in learning games object-centric representations (Delfosse et al., 2023b) in the first place.
 Impact statement
@@
 Even so, the inspectable nature of transparent approaches will allow to identify such potentially harmful misuses, or hidden misalignment.
+Acknowledgments
+This work has benefited from the HMWK projects ”The Third Wave of Artificial Intelligence - 3AI”, their joint support of the National Research Center for Applied Cybersecurity ATHENE, via the “SenPai: XReLeaS” project.
+It has also received support of Hessian.AI, as well as the Hessian research priority program LOEWE within the project WhiteBox, and the EU-funded “TANGO” project (EU Horizon 2023, GA No 57100431).
+The authors also would like to thank Stefan Lichtenstein, Raban Emunds for their help on refactoring the code, and Dwarak Vittal for his initial contributions.
 Appendix A Appendix
@@
 A.1 Atari games are most common set of environments
-We here show that the Atari games from the Arcade Learning Environment (Bellemare et al., 2012) is the most used benchmark to test reinforcement learning agents.
-Figure 5: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from (Delfosse et al., 2023a).
+Figure 6: The Atari Learning Environments is more used in scientific research than the next 88 other benchmarks together. Graph borrowed from [Delfosse et al., 2024b].
+We here show that the Atari games from the Arcade Learning Environment [Bellemare et al., 2012] is the most used benchmark to test reinforcement learning agents.
 A.2 Detailed numerical results
@@
 A detailed overview of hyperparameters and their corresponding values can be found in A.5.
-All agents are trained on gymnasium’s atari learning environments using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank).
+All agents are trained on gymnasium’s ALE using $8$ actors and $3$ training seeds ([0,16,32]+[0,16,32]+rank).
 Each training seed’s performance is evaluated every 500500k frames on $4$ differently seeded (42+42+training seed) environments for $8$ episodes each.
 After training, the best performing checkpoint is then ultimately evaluated on $4$ seeded ($123,456,789,1011$) test environments.
-The final return is determined by averaging the return over $5$ episodes per training-test seed pair and every training seed for the respective environment.
+The final return is determined by averaging the return over $5$ episodes per training-test seed and every training seed for the respective environment.
 We use deterministic actions for both evaluation and testing stages.
 Game
-SCoBots
-v5
-SCoBots
-guided-v5
-PPO
-ours-v5
-PPO
-ours-v4
-PPO
-Schulman et al.-v4
+SCoBots-v5
+NoGuidance
+SCoBots-v5
+PPO-v5
+ours
+PPO-v4
+ours
+PPO-v4
+Schulman et al.
 Random Human
-Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433.3\mbox{\scriptsize$\pm 2772.9$}$ $4532$ $210$ $8503$
+Asterix $5080\mbox{\scriptsize$\pm 614.8$}$ $5043.3\mbox{\scriptsize$\pm 729.9$}$ $2126.7\mbox{\scriptsize$\pm 148.6$}$ $10433\mbox{\scriptsize$\pm 2773$}$ $4532$ $210$ $8503$
 Bowling $106.6\mbox{\scriptsize$\pm 41.8$}$ $137.4\mbox{\scriptsize$\pm 24.1$}$ $102.2\mbox{\scriptsize$\pm 39.1$}$ $51.4\mbox{\scriptsize$\pm 18.0$}$ $40.1$ $23.1$ $160.7$
@@
 Neural refers to the agent that use a CNN instead of our ICB layers.
-We added the results reported in the original paper (Schulman et al., 2017) (original), which are on par with ours, as well as ones from a Random baseline and Humans (from (van Hasselt et al., 2016)).
+We added the results reported in the original paper Schulman et al.
+[2017] (original), which are on par with ours, as well as ones from a Random baseline and Humans (from van Hasselt et al.
+[2016]).
 Our agents have been trained with only 2020M frames.
@@
 A.3 Learning curves
-Figure 6: Training Return over frames ($\times 10^{6}$) seen for   guided SCoBots,   SCoBots and   neural agents.
+Figure 7: Training Return over frames ($\times 10^{6}$) seen for   SCoBots,   SCoBots w/o guidance and   neural agents.
 A.4 Agent reasoning through the decision trees
@@
 All Experiments were run on a AMD Ryzen 7 processor, 64GB of RAM and one NVIDIA GeForce RTX 2080 Ti GPU.
-Training a (i)SCoBots on $20\text{M}$ frames and using $8$ actors takes approximately $4$ hours.
+Training a SCoBot on $20\text{M}$ frames with $8$ actors takes approximately $8$ hours.
 We use the same PPO hyperparameters as the Schulman et al.
-(2017) agents that learned to master the games.
+[2017] agents that learned to master the games.
 For the Adam optimizer, SCoBots start with a slightly increased learning rate of $1\times 10^{-3}$ (compared to $2.5\times 10^{-4}$).
 The PPO implementation used and the respective MLP hyperparameters are based on stable-baselines3 Raffin et al.
-(2021).
+[2021].
 SCoBots have the same PPO hyperparameter values as deep agents but use MLPs ($2\times 64$) with ReLU activation functions as policy and value networks.
@@
 This is aligned with recommended best practices by Machado et al.
-(2018).
+[2018].
 The experiments using NoFrameskip-v4 utilize the same environment wrappers as OpenAI’s initial baselines implementation444github.com/openai/baselines.
@@
 Name Definition Description
-Properties class NAME object class (e.g. ”Agent”, ”Ball”, ”Ghost”)
+Properties class NAME object class (e.g. "Agent", "Ball", "Ghost")
 position $x,y$ position on the screen
@@
 The set $\mathcal{S}$ denotes the problem-specific set of raw environment states as defined by the RL problem, e.g., the space of RGB images $0,255]^{512\times 512\times 3}$.
-Similarly, $\mathcal{A}$ represents the action space, e.g., “move right” or “jump.”
+Similarly, $\mathcal{A}$ represents the action space, e.g., “move right" or “jump."
 Let $O_{0}$ be the set of all possible objects, identified by their ID and characterized by their properties like position, size, color, etc.
@@
 A.7 Pong misalignment problem
-The misalignment problem had previously been identified in other games such as the Coinrun platform game (Cobbe et al., 2019), in which an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time.
+The misalignment problem had previously been identified in other games such as the Coinrun platform game [Cobbe et al., 2019], in which an agent’s target-goal is to reach a coin which is always placed at the end of a level at training time.
 When the coin is repositioned at test time di Langosco et al.
@@
 We then compute different correlation coefficients, and obtain 99.6% as Pearson coefficient, 96.4% as Kendall coefficient and 99.5% as Spearman coefficient.
-We also collected importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from (Delfosse et al., 2024).
+We also collected importance maps of deep DQN RL agents playing Pong using ReLU and Rational activation functions (borrowed from Delfosse et al. [2024c].
 Figure 10: More importance maps of DQN agents playing Pong, with ReLU (left) and rational activation functions (right).
@@
 A.8.1 Ill-defined Objectives
-Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences (Henderson et al., 2017; Irving et al., 2018).
-Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common (Andrychowicz et al., 2017).
+Defining an accurate reward signal when creating an RL task is challenging, as it requires the designer to specify a precise and balanced reward function that effectively guides the agent towards the desired behavior while avoiding unintended consequences [Henderson et al., 2017, Irving et al., 2018].
+Shaping the reward after having observed (and understood) the non-intended behavior of the agent is very common [Andrychowicz et al., 2017].
 To realign our SCoBots agent, we provide them with a reward signal proportional to the progression to the joey:
@@
 A.8.2 Difficult Credit Assignment
-When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome (Mesnard et al., 2021).
+When an outcome is delayed or uncertain, it is challenging to identify the action that is relevant for that outcome [Mesnard et al., 2021].
 This problem arises in games like chess, where reward is passed to the agent only when the game is over.
@@
 This resulted in an agent that learned to stop itself in between the flags (depicted in Fig.
-11), showing once again, how difficult it is to create a reward that would favor a specific behavior.
+12), showing once again, how difficult it is to create a reward that would favor a specific behavior.
 To correct it, we therefore adjoined another signal to reward our SCoBot agents proportionally to the speed of the agent:
@@
 We observe that such agent do not rely on the enemy to master the game.
-Alternative techniques such as prioritized experience replay (Schaul et al., 2016) allow offline methods to learn in such environments, but providing a smoother reward signal is another elegant way to address sparsity.
+Alternative techniques such as prioritized experience replay [Schaul et al., 2016] allow offline methods to learn in such environments, but providing a smoother reward signal is another elegant way to address sparsity.
 Note that sparser environments exists (such as robotics ones).
@@
 Shortcut learning can be mitigated through the implementation of such reward signals that necessitate the genuine objective’s consideration.
+A.9 Computational load
+We here provide the overall computational walltime for training each agent type, on each Atari environment.
+When focusing on a limited number of objects, their computational time is nearly halved, particularly in environments with many objects.
+SCoBots Deep SCoBots (NG)
+Asterix $07:08$ $08:07$ $12:12$
+Bowling $07:17$ $11:13$ $08:20$
+Boxing $07:26$ $11:20$ $07:32$
+Freeway $07:34$ $11:13$ $09:44$
+Kangaroo $06:01$ $10:43$ $21:21$
+Pong $05:42$ $10:55$ $07:41$
+Skiing $07:02$ $10:37$ $08:38$
+Seaquest $08:53$ $10:51$ $25:12$
+Tennis $11:19$ $12:08$ $11:25$
+Mean $07:35$ $10:47$ $12:27$
+Max $11:19$ $12:08$ $25:12$
+Min $05:42$ $08:07$ $07:32$
+Table 5: SCoBots (with OCAtari) train faster than deep agents, particularly in environments with a limited number of objects.
+Computational training time of each method on each used environment (format $H$:$M$).
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: In this paper, we claim that object-centric RL agents that use successive concept bottlenecks can challenge deep RL ones, and that they allow for easy detection and corrections of diverse RL-specific caveats.
+This is shown in Figure 3, both for the performances, as well as for the interpretability and correction.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation, as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The paper discusses SCoBots’ limitations in the Limitation section included in the main paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: The paper does not include theoretical results.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: The paper presents the experimental setup and evaluation process in detail, both in the main paper and appendix.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24.
+Justification: The paper includes all information necessary to reproduce the experimental results and provides open access to a git code repository (using anonymous-github until publication, then the code will be publicly available on GitHub).
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The experimental setup is described in the main paper, all hyperparameter values are listed in detail in the appendix and the code is available.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: Error bars are reported in the main paper, detailed numerical results included in the appendix.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The compute resources and time required are listed in detail in Appendix A5. The total compute time for our experiments was approximately $1200$ CPU+GPU hours.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44.
+Justification: We discuss the Potential Harms and negative societal impacts that interpretable algorithms can provide, as a misintended person with discriminative behavior could increase the misalignment of such agents.
+We hope that overall, the transparency of agents’ decisions will result in positive societal impact.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: The paper discusses societal impacts in an impact statement included in the main paper.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We simply train transparent agents on Atari environments, we do not see societal negative impact that these agents could bring.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite every creator of the used environments and RL training repositories, all of which are publicly available and freely distributed.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: Our code is commented, but this does not constitute an asset.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.